### v1

# On solutions of reduced biquaternion equality constrained least squares problem and their relative forward error bound 22footnotemark: 2

Abstract
This study focuses on addressing the challenge of solving the reduced biquaternion equality constrained least squares (RBLSE) problem. We develop algebraic techniques to derive both complex and real solutions for the RBLSE problem by utilizing the complex and real forms of reduced biquaternion matrices. Additionally, we conduct a perturbation analysis for the RBLSE problem and establish an upper bound for the relative forward error of these solutions. Numerical examples are presented to illustrate the effectiveness of the proposed approaches and to verify the accuracy of the established upper bound for the relative forward errors.
Keywords. Real representation matrix, Reduced biquaternion, Complex representation matrix, Equality constrained least squares problem.
AMS subject classification. 15A60, 15B33, 65F20, 65H10.
1 Introduction
Reduced biquaternions, first introduced by Segre in $1892$, have become fundamental in various domains such as signal processing, image analysis, and neural networks [4, 5, 6, 7]. Unlike standard quaternions, reduced biquaternions, often referred to as commutative quaternions, possess the unique property of commutative multiplication. This property not only simplifies mathematical operations but also significantly boosts computational efficiency, as demonstrated in [9, 10].
The wide-ranging applications of reduced biquaternions have spurred significant research into both the theoretical and numerical aspects of matrices and matrix equations involving commutative quaternions. For instance, algorithms for calculating eigenvalues, eigenvectors, and singular value decompositions of reduced biquaternion matrices are detailed in [10]. Zhang et al. explored the singular value decomposition and generalized inverse of these matrices in [14], and further investigated the diagonalization process in [15], where they established the necessary and sufficient conditions for diagonalization and introduced two numerical methods to facilitate this task. In [3], authors discussed the LU decomposition of reduced biquaternion matrices. In [8], Guo et al. explored eigen-problems of reduced biquaternion matrices, while in [11], Yuan et al. discussed the Hermitian solution of the reduced biquaternion matrix equation $\left(AXB,CXD\right)=\left(E,F\right)$. In [2], Ding et al. explored the solutions to matrix equations $XF-AX=BY$ and $XF-A\widetilde{X}=BY$ over commutative quaternions.
In many practical applications, determining the solution to linear system, typically expressed as $AX\approx B$, is a common challenge. The least squares method is a well-known approach for addressing this problem. Prior research, such as the work in [12], has explored least squares solutions for matrix equations like $AX=B$ and $AXC=B$ over commutative quaternions. The equality constrained least squares problem is a further study of the least squares problem, and in [13], the authors discussed solution techniques for computing the reduced biquaternion solution of the reduced biquaternion equality constrained least squares (RBLSE) problem. In this manuscript, we advance the research by developing methods to obtain both complex and real solutions for the RBLSE problem. Additionally, we establish an upper bound for the relative forward error associated with these solutions, ensuring the accuracy and reliability of our method in addressing RBLSE problems.
The organization of the manuscript is as follows: Section 2 covers the notation and preliminary concepts. Section 3 outlines the method for finding complex solution of the RBLSE problem, while Section 4 focuses on the technique for finding real solutions to the same problem. Lastly, Section 5 presents the numerical validation of the proposed methods.
2 Notation and preliminaries
2.1 Notation
In this paper, we adopt the following notations: $\mathbb{Q}_{\mathbb{R}}^{m\times n}$, ${\mathbb{C}}^{m\times n}$, and ${\mathbb{R}}^{m\times n}$ represent the sets of all $m\times n$ reduced biquaternion, complex, and real matrices. The notations $\left\lVert\cdot\right\rVert_{2}$ and $\left\lVert\cdot\right\rVert_{F}$ refer to the $2$-norm and Frobenius norm, respectively. For matrices $A\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$ and $B\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $\left[A,B\right]$ represents their concatenation, i.e., $\begin{bmatrix}A&B\end{bmatrix}\in\mathbb{Q}_{\mathbb{R}}^{m\times(n+d)}$. For a matrix $A\in{\mathbb{C}}^{m\times n}$, the symbols $A^{T}$, $A^{H}$, and $A^{\dagger}$ correspond to the transpose, conjugate transpose, and Moore-Penrose inverse of $A$, respectively. Additionally, $A^{-1}$ indicates the inverse of $A$ if it exists.
The Matlab functions $randn(m,n)$ and $rand(m,n)$ produce $m\times n$ matrices filled with normally distributed and uniformly distributed random values, respectively. The abbreviations RBLSE (reduced biquaternion equality constrained least squares) and LSE (equality constrained least squares) will be used throughout this paper.
2.2 Preliminaries
A reduced biquaternion is defined as $\zeta=p_{0}+p_{1}\textit{{i}}+p_{2}\textit{{j}}+p_{3}\textit{{k}}$, which can also be expressed as $\zeta=\left(p_{0}+p_{1}\textit{{i}}\right)+\left(p_{2}+p_{3}\textit{{i}}\right)\textit{{j}}=r_{1}+r_{2}\textit{{j}}$. Here, $p_{i}\in{\mathbb{R}}$ for $i=0,1,2,3$ and $r_{i}\in{\mathbb{C}}$ for $i=1,2$. The defining properties include $\textit{{i}}^{2}=\textit{{k}}^{2}=-1$ and $\textit{{j}}^{2}=1$, with multiplication rules $\textit{{j}}=\textit{{j}}\textit{{i}}=\textit{{k}},\;\textit{{j}}\textit{{k}}=\textit{{k}}\textit{{j}}=\textit{{i}},$ i , and $\textit{{k}}\textit{{i}}=\textit{{i}}\textit{{k}}=-\textit{{j}}$. The norm of $\zeta$ is computed as $\left\lVert\zeta\right\rVert=\sqrt{p_{0}^{2}+p_{1}^{2}+p_{2}^{2}+p_{3}^{2}}=\sqrt{|r_{1}|^{2}+|r_{2}|^{2}}$. For a matrix $M=\left(m_{ij}\right)\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, the Frobenius norm [11] is given by:
$\left\lVert M\right\rVert_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}\left\lVert m_{ij}\right\rVert^{2}}.$ (1)
Suppose $M=M_{0}+M_{1}\textit{{i}}+M_{2}\textit{{j}}+M_{3}\textit{{k}}=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, where $M_{i}\in{\mathbb{R}}^{m\times n}$ for $i=0,1,2,3$ and $N_{i}\in{\mathbb{C}}^{m\times n}$ for $i=1,2$. The real and complex representations of matrix $M$, denoted as $M^{R}$ and $M^{C}$, are given by:
$M^{R}=\begin{bmatrix}M_{0}&-M_{1}&M_{2}&-M_{3}\\ M_{1}&M_{0}&M_{3}&M_{2}\\ M_{2}&-M_{3}&M_{0}&-M_{1}\\ M_{3}&M_{2}&M_{1}&M_{0}\end{bmatrix},\;\;M^{C}=\begin{bmatrix}N_{1}&N_{2}\\ N_{2}&N_{1}\end{bmatrix}.$ (2)
Let $M_{c}^{R}$ represent the first block column of the matrix $M^{R}$. More precisely, $M_{c}^{R}$ is defined as: $M^{R}_{c}=\left[M_{0}^{T},M_{1}^{T},M_{2}^{T},M_{3}^{T}\right]^{T}.$ The matrix $M^{R}$ can be represented using $M^{R}_{c}$ as:
$M^{R}=\left[M_{c}^{R},Q_{m}M_{c}^{R},R_{m}M_{c}^{R},S_{m}M_{c}^{R}\right],$ (3)
where
$Q_{m}=\begin{bmatrix}0&-I_{m}&0&0\\ I_{m}&0&0&0\\ 0&0&0&-I_{m}\\ 0&0&I_{m}&0\end{bmatrix},\;R_{m}=\begin{bmatrix}0&0&I_{m}&0\\ 0&0&0&I_{m}\\ I_{m}&0&0&0\\ 0&I_{m}&0&0\end{bmatrix},\mbox{and}\;S_{m}=\begin{bmatrix}0&0&0&-I_{m}\\ 0&0&I_{m}&0\\ 0&-I_{m}&0&0\\ I_{m}&0&0&0\end{bmatrix}.$ (4)
We have
$\left\lVert M\right\rVert_{F}=\frac{1}{2}\left\lVert M^{R}\right\rVert_{F}=\left\lVert M^{R}_{c}\right\rVert_{F}.$ (5)
Lemma 2.1.
Let $P\in{\mathbb{R}}^{m\times n}$, $Q\in{\mathbb{R}}^{m\times d}$, $R\in{\mathbb{R}}^{m\times p}$, and $S\in{\mathbb{R}}^{m\times q}$. If $\|P\|_{F}=\|Q\|_{F}=\|R\|_{F}=\|S\|_{F}$, then we have
$\left\lVert P\right\rVert_{F}=\frac{1}{2}\left\lVert\left[P,Q,R,S\right]\right\rVert_{F}.$
Let $M_{c}^{C}$ represent the first block column of the matrix $M^{C}$. More precisely, $M_{c}^{C}$ is defined as: $M^{C}_{c}=\left[M_{1}^{T},M_{2}^{T}\right]^{T}.$ The matrix $M^{C}$ can be represented using $M^{C}_{c}$ as:
$M^{C}=\left[M_{c}^{C},P_{m}M_{c}^{C}\right],$ (6)
where
$P_{m}=\begin{bmatrix}0&I_{m}\\ I_{m}&0\end{bmatrix}.$ (7)
We have
$\left\lVert M\right\rVert_{F}=\frac{1}{\sqrt{2}}\left\lVert M^{C}\right\rVert_{F}=\left\lVert M^{C}_{c}\right\rVert_{F}.$ (8)
Lemma 2.2.
Let $P\in{\mathbb{C}}^{m\times n}$ and $Q\in{\mathbb{C}}^{m\times d}$. If $\|P\|_{F}=\|Q\|_{F}$, then we have
$\left\lVert P\right\rVert_{F}=\frac{1}{\sqrt{2}}\left\lVert\left[P,Q\right]\right\rVert_{F}.$
Lemma 2.3.
For $\alpha\in{\mathbb{R}}$, $\beta\in{\mathbb{C}}$, $P,Q\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, and $R\in\mathbb{Q}_{\mathbb{R}}^{n\times t}$, the following relationships hold:
  1. $(a)$
    $P=Q\iff P^{C}=Q^{C}\iff P^{R}=Q^{R}$.
  2. $(b)$
    $\left(P+Q\right)^{R}=P^{R}+Q^{R}$, $\left(P+Q\right)^{C}=P^{C}+Q^{C}$.
  3. $(c)$
    $\left(\alpha P\right)^{R}=\alpha P^{R}$, $\left(\beta P\right)^{C}=\beta P^{C}$.
  4. $(d)$
    $\left(PR\right)^{R}=P^{R}R^{R}$, $\left(PR\right)^{C}=P^{C}R^{C}$.
3 An algebraic technique for complex solution of RBLSE problem
This section focuses on an algebraic approach to derive the complex solution for the RBLSE problem. The method is based on analyzing the solution of the associated complex LSE problem. Suppose
$A$ $=$ $M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\;\;\;\;B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (9)
$C$ $=$ $R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\;\;\;\;D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (10)
We will limit our discussion to the scenario where $m\geq n+d$, and both matrices $C$ and $C^{C}_{c}$ have full row rank. The RBLSE problem can then be stated as follows:
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;CX=D.$ (11)
Consider a complex LSE problem
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{C}_{c}X=D^{C}_{c}.$ (12)
Let the QR factorization of $\left(C^{C}_{c}\right)^{H}$ be as follows:
$\left(C^{C}_{c}\right)^{H}=\widetilde{Q}\begin{bmatrix}\widetilde{R}\\ 0\end{bmatrix},$ (13)
where $\widetilde{Q}\in{\mathbb{C}}^{n\times n}$ is a unitary matrix and $\widetilde{R}\in{\mathbb{C}}^{2p\times 2p}$ is a nonsingular upper triangular matrix. Now, partition $A^{C}_{c}\widetilde{Q}$ as:
$A^{C}_{c}\widetilde{Q}=\left[\widetilde{P}_{1},\widetilde{P}_{2}\right],$ (14)
where $\widetilde{P}_{1}\in{\mathbb{C}}^{2m\times 2p}$ and $\widetilde{P}_{2}\in{\mathbb{C}}^{2m\times(n-2p)}$.
Theorem 3.1.
Consider the RBLSE problem outlined in (11) and the complex LSE problem in (12), with notations specified in (13) and (14). For a matrix $X\in{\mathbb{C}}^{n\times d}$, $X$ is a complex solution of the RBLSE problem (11) if and only if $X$ solves the complex LSE problem (12). In this scenario, the unique solution $X$ with minimum norm can be expressed as:
$X=\widetilde{Q}\begin{bmatrix}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}\\ \widetilde{P}_{2}^{\dagger}\left(B^{C}_{c}-\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}\right)\end{bmatrix}.$ (15)
Proof.
If $X\in{\mathbb{C}}^{n\times d}$ is a solution of the complex LSE problem (12), then
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\textrm{min},\;\;C^{C}_{c}X=D^{C}_{c}.$ (16)
The Frobenius norm of a complex matrix remains invariant under unitary transformations. Since the matrix $P_{m}$ in (7) is unitary, it follows that:
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\left\lVert P_{m}\left(A^{C}_{c}X-B^{C}_{c}\right)\right\rVert_{F}.$
Using equations (2), (6), (8), along with Lemmas 2.2 and 2.3, we obtain
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}$ $=$ $\frac{1}{\sqrt{2}}\left\lVert\left[\left(A^{C}_{c}X-B^{C}_{c}\right),P_{m}\left(A^{C}_{c}X-B^{C}_{c}\right)\right]\right\rVert_{F}$
$=$ $\frac{1}{\sqrt{2}}\left\lVert\left[A^{C}_{c}X,P_{m}A^{C}_{c}X\right]-\left[B^{C}_{c},P_{m}B^{C}_{c}\right]\right\rVert_{F}$
$=$ $\frac{1}{\sqrt{2}}\left\lVert\left[A^{C}_{c},P_{m}A^{C}_{c}\right]\begin{bmatrix}X&0\\ 0&X\end{bmatrix}-\left[B^{C}_{c},P_{m}B^{C}_{c}\right]\right\rVert_{F}$
$=$ $\frac{1}{\sqrt{2}}\left\lVert A^{C}X^{C}-B^{C}\right\rVert_{F}$
$=$ $\frac{1}{\sqrt{2}}\left\lVert\left(AX-B\right)^{C}\right\rVert_{F}$
$=$ $\left\lVert AX-B\right\rVert_{F}.$
From (16), we obtain
$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\left\lVert AX-B\right\rVert_{F}=\textrm{min},$ (17)
and
$\left[C^{C}_{c},P_{p}C^{C}_{c}\right]\begin{bmatrix}X&0\\ 0&X\end{bmatrix}=\left[D^{C}_{c},P_{p}D^{C}_{c}\right].$
Using (6), we know that $C^{C}=\left[C^{C}_{c},P_{p}C^{C}_{c}\right]$ and $D^{C}=\left[D^{C}_{c},P_{p}D^{C}_{c}\right]$. Applying this, we get
$C^{C}X^{C}$ $=$ $D^{C},$
$\left(CX\right)^{C}$ $=$ $D^{C},$
$CX$ $=$ $D.$ (18)
By combining (17) and (18), we conclude that $X\in{\mathbb{C}}^{n\times d}$ is a complex solution to the RBLSE problem (11), and vice versa.
To find the expression for $X$, we solve the complex LSE problem (12). Set
$\widetilde{Q}^{H}X=\begin{bmatrix}Y\\ Z\end{bmatrix},$
where $Y\in{\mathbb{C}}^{2p\times d}$ and $Z\in{\mathbb{C}}^{(n-2p)\times d}$. Equation 12 can be rewritten as
$\min_{X}\left\lVert A^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X-B^{C}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X=D^{C}_{c}.$ (19)
Utilizing (13), we have
$C^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X=D^{C}_{c}\implies\begin{bmatrix}\widetilde{R}^{H}&0\end{bmatrix}\begin{bmatrix}Y\\ Z\end{bmatrix}=D^{C}_{c}.$
Since $\widetilde{R}^{H}$ is a nonsingular matrix, we get $Y=\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$. Using (14), Equation 19 takes the form
$\min_{Z}\left\lVert\widetilde{P}_{2}Z-\left(B^{C}_{c}-\widetilde{P}_{1}Y\right)\right\rVert_{F}.$
The minimum norm solution of the above least squares problem is $Z=\widetilde{P}_{2}^{\dagger}\left(B^{C}_{c}-\widetilde{P}_{1}Y\right)$. Thus, we can derive the desired expression for $X$. ∎
Next, we aim to examine how perturbations in $A$, $B$, $C$, and $D$ affect the complex solution $X_{CL}$ of the RBLSE problem (11). Let
$\widehat{A}=A+\Delta A,\;\widehat{B}=B+\Delta B,\;\widehat{C}=C+\Delta C,\;\textrm{and}\;\widehat{D}=D+\Delta D,$ (20)
where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively. We assume that the perturbation $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to guarantee that the perturbed matrices $\widehat{C}$ and $\widehat{C}^{C}_{c}$ retain full row rank. These perturbations are measured normwise by the smallest $\epsilon$ for which
$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\left\lVert\Delta C\right\rVert_{F}\leq\epsilon\left\lVert C\right\rVert_{F},\;\textrm{and}\;\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (21)
Let $\widehat{X}_{CL}$ be the complex solution to the perturbed RBLSE problem
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\;\;\textrm{subject to}\;\;\widehat{C}X=\widehat{D},$ (22)
and let $\Delta X_{CL}=\widehat{X}_{CL}-X_{CL}$.
Theorem 3.2.
Consider the RBLSE problem outlined in (11) and the perturbed RBLSE problem described in (22). If the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are sufficiently small, as described in (21), then we have
$\begin{split}\frac{\left\lVert\Delta X_{CL}\right\rVert_{F}}{\left\lVert X_{CL}\right\rVert_{F}}\leq\epsilon&\left(\mathcal{K}_{A}^{C}\left(\frac{\left\lVert D^{C}_{c}\right\rVert_{F}}{\left\lVert C^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}+1\right)+\mathcal{K}_{B}^{C}\left(\frac{\left\lVert B^{C}_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}+1\right)\right.\\ &\left.+\left(\mathcal{K}_{B}^{C}\right)^{2}\left(\frac{\left\lVert C^{C}_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}}\left\lVert A^{C}_{c}\mathcal{L}_{c}\right\rVert_{2}+1\right)\frac{\left\lVert R_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}\right)+O(\epsilon^{2})\equiv U_{CL},\end{split}$ (23)
where
$\mathcal{K}_{B}^{C}$ $=$ $\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert\left(A^{C}_{c}P_{c}\right)^{\dagger}\right\rVert_{2},\;\;\;\mathcal{K}_{A}^{C}=\left\lVert C^{C}_{c}\right\rVert_{F}\left\lVert\mathcal{L}_{c}\right\rVert_{2},\;\;\;\mathcal{L}_{c}=\left(I_{n}-\left(A^{C}_{c}P_{c}\right)^{\dagger}A^{C}_{c}\right)\left(C^{C}_{c}\right)^{\dagger},\;\;\;\;$
$P_{c}$ $=$ $I_{n}-\left(C^{C}_{c}\right)^{\dagger}C^{C}_{c},\;\;\;\;\;\;\;\;\;\;\;R_{c}=B^{C}_{c}-A^{C}_{c}X_{CL}.$
Proof.
The perturbed complex LSE problem corresponding to the perturbed RBLSE problem (22) is given by:
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{C}_{c}X-\left(\widehat{B}\right)^{C}_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(\widehat{C}\right)^{C}_{c}X=\left(\widehat{D}\right)^{C}_{c}.$ (24)
Using Theorem 3.1, we know that $\widehat{X}_{CL}$ is the solution to the perturbed complex LSE problem (24). From (20) and utilizing Lemma 2.3, we have
$\left(\widehat{A}\right)^{C}_{c}=A^{C}_{c}+\left(\Delta A\right)^{C}_{c},\;\left(\widehat{B}\right)^{C}_{c}=B^{C}_{c}+\left(\Delta B\right)^{C}_{c},\;\left(\widehat{C}\right)^{C}_{c}=C^{C}_{c}+\left(\Delta C\right)^{C}_{c},\;\textrm{and}\;\left(\widehat{D}\right)^{C}_{c}=D^{C}_{c}+\left(\Delta D\right)^{C}_{c}.$
Thus, the perturbed complex LSE problem (24) can be rewritten as:
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(A^{C}_{c}+\left(\Delta A\right)^{C}_{c}\right)X-\left(B^{C}_{c}+\left(\Delta B\right)^{C}_{c}\right)\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(C^{C}_{c}+\left(\Delta C\right)^{C}_{c}\right)X=\left(D^{C}_{c}+\left(\Delta D\right)^{C}_{c}\right).$ (25)
Using (8) and (21), we can establish the following bounds for the perturbation:
$\left\lVert\left(\Delta A\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{C}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta B\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{C}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta C\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{C}_{c}\right\rVert_{F},\;\textrm{and}\;\left\lVert\left(\Delta D\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{C}_{c}\right\rVert_{F}.$ (26)
With the perturbed problem (25) and the bounds in (26), and using Theorem 3.1, the sensitivity analysis of the complex solution to the RBLSE problem (11) reduces to evaluating the sensitivity of the solution to the complex LSE problem (12). Consequently, the upper bound $U_{CL}$ for the relative forward error of the complex solution to the RBLSE problem can be obtained from [1, Equation $4.11$]. ∎
4 An algebraic technique for real solution of RBLSE problem
This section focuses on an algebraic approach to derive the real solution for the RBLSE problem. The method is based on analyzing the solution of the associated real LSE problem. Suppose
$A$ $=$ $A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\;\;\;\;B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (27)
$C$ $=$ $C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\;\;\;\;\;D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (28)
We will limit our discussion to the scenario where $m\geq n+d$, and both matrices $C$ and $C^{R}_{c}$ have full row rank. The RBLSE problem can then be stated as follows:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;CX=D.$ (29)
Consider a real LSE problem
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{R}_{c}X=D^{R}_{c}.$ (30)
Let the QR factorization of $\left(C^{R}_{c}\right)^{T}$ be as follows:
$\left(C^{R}_{c}\right)^{T}=\bar{Q}\begin{bmatrix}\bar{R}\\ 0\end{bmatrix},$ (31)
where $\bar{Q}\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix and $\bar{R}\in{\mathbb{R}}^{4p\times 4p}$ is a nonsingular upper triangular matrix. Now, partition $A^{R}_{c}\bar{Q}$ as:
$A^{R}_{c}\bar{Q}=\left[\bar{P}_{1},\bar{P}_{2}\right],$ (32)
where $\bar{P}_{1}\in{\mathbb{R}}^{4m\times 4p}$ and $\bar{P}_{2}\in{\mathbb{R}}^{4m\times(n-4p)}$.
Theorem 4.1.
Consider the RBLSE problem outlined in (29) and the real LSE problem in (30), with notations specified in (31) and (32). For a matrix $X\in{\mathbb{R}}^{n\times d}$, $X$ is a real solution of the RBLSE problem (29) if and only if $X$ solves the real LSE problem (30). In this scenario, the unique solution with minimum norm $X$ can be expressed as:
$X=\bar{Q}\begin{bmatrix}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}\\ \bar{P}_{2}^{\dagger}\left(B^{R}_{c}-\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}\right)\end{bmatrix}.$ (33)
Proof.
If $X\in{\mathbb{R}}^{n\times d}$ is a solution of the real LSE problem (30), then
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\textrm{min},\;\;C^{R}_{c}X=D^{R}_{c}.$ (34)
The Frobenius norm of a real matrix remains invariant under orthogonal transformations. Since the matrices $Q_{m}$, $R_{m}$, and $S_{m}$ in (4) are orthogonal, it follows that:
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\left\lVert Q_{m}\left(A^{R}_{c}X-B^{R}_{c}\right)\right\rVert_{F}=\left\lVert R_{m}\left(A^{R}_{c}X-B^{R}_{c}\right)\right\rVert_{F}=\left\lVert S_{m}\left(A^{R}_{c}X-B^{R}_{c}\right)\right\rVert_{F}.$
Using equations (2), (3), (5), along with Lemmas 2.1 and 2.3, we obtain
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}$ $=$ $\frac{1}{2}\left\lVert\left[\left(A^{R}_{c}X-B^{R}_{c}\right),Q_{m}\left(A^{R}_{c}X-B^{R}_{c}\right),R_{m}\left(A^{R}_{c}X-B^{R}_{c}\right),S_{m}\left(A^{R}_{c}X-B^{R}_{c}\right)\right]\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert\left[A^{R}_{c}X,Q_{m}A^{R}_{c}X,R_{m}A^{R}_{c}X,S_{m}A^{R}_{c}X\right]-\left[B^{R}_{c},Q_{m}B^{R}_{c},R_{m}B^{R}_{c},S_{m}B^{R}_{c}\right]\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert\left[A^{R}_{c},Q_{m}A^{R}_{c},R_{m}A^{R}_{c},S_{m}A^{R}_{c}\right]\begin{bmatrix}X&0&0&0\\ 0&X&0&0\\ 0&0&X&0\\ 0&0&0&X\end{bmatrix}-\left[B^{R}_{c},Q_{m}B^{R}_{c},R_{m}B^{R}_{c},S_{m}B^{R}_{c}\right]\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert A^{R}X^{R}-B^{R}\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert\left(AX-B\right)^{R}\right\rVert_{F}$
$=$ $\left\lVert AX-B\right\rVert_{F}.$
From (34), we obtain
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\left\lVert AX-B\right\rVert_{F}=\textrm{min},$ (35)
and
$\left[C^{R}_{c},Q_{p}C^{R}_{c},R_{p}C^{R}_{c},S_{p}C^{R}_{c}\right]\begin{bmatrix}X&0&0&0\\ 0&X&0&0\\ 0&0&X&0\\ 0&0&0&X\end{bmatrix}=\left[D^{R}_{c},Q_{p}D^{R}_{c},R_{p}D^{R}_{c},S_{p}D^{R}_{c}\right].$
Using (3), we know that $C^{R}=\left[C^{R}_{c},Q_{p}C^{R}_{c},R_{p}C^{R}_{c},S_{p}C^{R}_{c}\right]$ and $D^{R}=\left[D^{R}_{c},Q_{p}D^{R}_{c},R_{p}D^{R}_{c},S_{p}D^{R}_{c}\right]$. Applying this, we get
$C^{R}X^{R}$ $=$ $D^{R},$
$(CX)^{R}$ $=$ $D^{R},$
$CX$ $=$ $D.$ (36)
By combining (35) and (36), we conclude that $X\in{\mathbb{R}}^{n\times d}$ is a real solution to the RBLSE problem (29), and vice versa. The expression for $X$ can be obtained by following the proof method of Theorem 3.1. ∎
Next, we aim to examine how perturbations in $A$, $B$, $C$, and $D$ affect the real solution $X_{RL}$ of the RBLSE problem (29). Let
$\widehat{A}=A+\Delta A,\;\widehat{B}=B+\Delta B,\;\widehat{C}=C+\Delta C,\;\textrm{and}\;\widehat{D}=D+\Delta D,$ (37)
where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively. We assume that the perturbation $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to ensure that the perturbed matrices $\widehat{C}$ and $\widehat{C}^{R}_{c}$ retain full row rank. These perturbations are measured normwise by the smallest $\epsilon$ for which
$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\textrm{and}\;\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (38)
Let $\widehat{X}_{RL}$ be the real solution to the perturbed RBLSE problem
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\;\;\textrm{subject to}\;\;\widehat{C}X=\widehat{D},$ (39)
and let $\Delta X_{RL}=\widehat{X}_{RL}-X_{RL}$.
Theorem 4.2.
Consider the RBLSE problem outlined in (29) and the perturbed RBLSE problem described in (39). If the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are sufficiently small, as described in (38), then we have
$\begin{split}\frac{\left\lVert\Delta X_{RL}\right\rVert_{F}}{\left\lVert X_{RL}\right\rVert_{F}}\leq\epsilon&\left(\mathcal{K}_{A}^{R}\left(\frac{\left\lVert D^{R}_{c}\right\rVert_{F}}{\left\lVert C^{R}_{c}\right\rVert_{F}\left\lVert X_{RL}\right\rVert_{F}}+1\right)+\mathcal{K}_{B}^{R}\left(\frac{\left\lVert B^{R}_{c}\right\rVert_{F}}{\left\lVert A^{R}_{c}\right\rVert_{F}\left\lVert X_{RL}\right\rVert_{F}}+1\right)\right.\\ &\left.+\left(\mathcal{K}_{B}^{R}\right)^{2}\left(\frac{\left\lVert C^{R}_{c}\right\rVert_{F}}{\left\lVert A^{R}_{c}\right\rVert_{F}}\left\lVert A^{R}_{c}\mathcal{L}_{r}\right\rVert_{2}+1\right)\frac{\left\lVert R_{r}\right\rVert_{F}}{\left\lVert A^{R}_{c}\right\rVert_{F}\left\lVert X_{RL}\right\rVert_{F}}\right)+O(\epsilon^{2})\equiv U_{RL},\end{split}$ (40)
where
$\mathcal{K}_{B}^{R}$ $=$ $\left\lVert A^{R}_{c}\right\rVert_{F}\left\lVert\left(A^{R}_{c}P_{r}\right)^{\dagger}\right\rVert_{2},\;\;\;\mathcal{K}_{A}^{R}=\left\lVert C^{R}_{c}\right\rVert_{F}\left\lVert\mathcal{L}_{r}\right\rVert_{2},\;\;\;\mathcal{L}_{r}=\left(I_{n}-\left(A^{R}_{c}P_{r}\right)^{\dagger}A^{R}_{c}\right)\left(C^{R}_{c}\right)^{\dagger},\;\;\;\;$
$P_{r}$ $=$ $I_{n}-\left(C^{R}_{c}\right)^{\dagger}C^{R}_{c},\;\;\;\;\;\;\;\;\;\;\;R_{r}=B^{R}_{c}-A^{R}_{c}X_{RL}.$
Proof.
The perturbed real LSE problem corresponding to the perturbed RBLSE problem (39) is given by:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{R}_{c}X-\left(\widehat{B}\right)^{R}_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(\widehat{C}\right)^{R}_{c}X=\left(\widehat{D}\right)^{R}_{c}.$ (41)
Using Theorem 4.1, we know that $\widehat{X}_{RL}$ is the solution to the perturbed real LSE problem (41). From (37) and utilizing Lemma 2.3, we have
$\left(\widehat{A}\right)^{R}_{c}=A^{R}_{c}+\left(\Delta A\right)^{R}_{c},\;\left(\widehat{B}\right)^{R}_{c}=B^{R}_{c}+\left(\Delta B\right)^{R}_{c},\;\left(\widehat{C}\right)^{R}_{c}=C^{R}_{c}+\left(\Delta C\right)^{R}_{c},\;\textrm{and}\;\left(\widehat{D}\right)^{R}_{c}=D^{R}_{c}+\left(\Delta D\right)^{R}_{c}.$
Thus, the perturbed real LSE problem (41) can be rewritten as:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(A^{R}_{c}+\left(\Delta A\right)^{R}_{c}\right)X-\left(B^{R}_{c}+\left(\Delta B\right)^{R}_{c}\right)\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(C^{R}_{c}+\left(\Delta C\right)^{R}_{c}\right)X=\left(D^{R}_{c}+\left(\Delta D\right)^{R}_{c}\right).$ (42)
Using (5) and (38), we can establish the following bounds for the perturbation:
$\left\lVert\left(\Delta A\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{R}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta B\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{R}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta C\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{R}_{c}\right\rVert_{F},\;\textrm{and}\;\left\lVert\left(\Delta D\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{R}_{c}\right\rVert_{F}.$ (43)
With the perturbed problem (42) and the bounds in (43), and using Theorem 4.1, the sensitivity analysis of the real solution to the RBLSE problem (29) reduces to evaluating the sensitivity of the solution to the real LSE problem (30). Consequently, the upper bound $U_{RL}$ for the relative forward error of the real solution to the RBLSE problem can be obtained from [1, Equation $4.11$]. ∎
5 Numerical Verification
Building on the previous discussions, this section presents numerical algorithms designed to find special solutions to the RBLSE problem. We also include numerical examples to validate these algorithms. Additionally, we assess the upper bound for the relative forward error of both complex and real solutions to the RBLSE problem.
Algorithm 1 For the complex solution of the RBLSE problem (11)
Input: $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that both matrices $C$ and $C^{C}_{c}$ have full row rank.
Output: $X$.
  1. 1.
    Find the QR factorization of $\left(C^{C}_{c}\right)^{H}$, as described in (13).
  2. 2.
    Partition the matrix $A^{C}_{c}\widetilde{Q}$ as shown in (14).
  3. 3.
    The complex solution $X$ to the RBLSE problem (11) is given by (15).
Algorithm 2 For the real solution of the RBLSE problem (29)
Input: $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that both matrices $C$ and $C^{R}_{c}$ have full row rank.
Output: $X$.
  1. 1.
    Find the QR factorization of $\left(C^{R}_{c}\right)^{T}$, as described in (31).
  2. 2.
    Partition the matrix $A^{R}_{c}\bar{Q}$ as shown in (32).
  3. 3.
    The real solution $X$ to the RBLSE problem (29) is given by (33).
Next, we provide examples to evaluate the effectiveness of the proposed algorithms. All computations were performed using a computer equipped with an Intel Core $i7-9700$ processor at $3.00$ GHz and $16$ GB of RAM, running MATLAB $R2021b$ software.
Example 5.1.
Let $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}$. Let
$M_{i}$ $=$ $rand(m,n)+rand(m,n)\textit{{i}}\in{\mathbb{C}}^{m\times n},\;\textrm{for}\;i=1,2,$
$N_{i}$ $=$ $rand(m,d)+rand(m,d)\textit{{i}}\in{\mathbb{C}}^{m\times d},\;\textrm{for}\;i=1,2,$
$R_{i}$ $=$ $rand(p,n)+rand(p,n)\textit{{i}}\in{\mathbb{C}}^{p\times n},\;\textrm{for}\;i=1,2,$
$S_{i}$ $=$ $rand(p,d)+rand(p,d)\textit{{i}}\in{\mathbb{C}}^{p\times d},\;\textrm{for}\;i=1,2.$
Take $m=40t$, $n=6t$, $p=2t$, and $d=3$. Here, $t$ is an arbitrary number. We apply Algorithm 1 to determine the complex solution of the RBLSE problem. Let $X_{CL}$ be the complex solution of the RBLSE problem
$\min_{X,R_{c}}\left\lVert R_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;AX=B+R_{c},\;CX=D.$
Let $\epsilon_{1}=\left\lVert AX_{CL}-(B+R_{c})\right\rVert_{F}$ and $\epsilon_{2}=\left\lVert CX_{CL}-D\right\rVert_{F}$. In Table 1, we compute $\epsilon_{1}$ and $\epsilon_{2}$ for different values of $t$.
$t$ $\epsilon_{1}$ $\epsilon_{2}$
$1$ $8.5131\times 10^{-16}$ $2.0907\times 10^{-15}$
$3$ $7.0956\times 10^{-16}$ $2.5624\times 10^{-15}$
$5$ $1.0499\times 10^{-15}$ $3.7683\times 10^{-15}$
$7$ $1.2804\times 10^{-15}$ $4.5681\times 10^{-15}$
$9$ $1.4232\times 10^{-15}$ $8.1546\times 10^{-15}$
Table 1: Computational accuracy of Algorithm 1 for computing the complex solution of the RBLSE problem (11)
Table 1 shows that the errors $\epsilon_{1}$ and $\epsilon_{2}$ across different values of $t$ are consistently below $10^{-15}$. This indicates that Algorithm 1 is highly effective in determining the complex solution for the RBLSE problem (11).
Example 5.2.
In Example 5.1, we introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively, to examine how these small perturbations affect the complex solution $X_{CL}$ of the RBLSE problem (11). These perturbations are measured normwise by the smallest $\epsilon$, as in (21). We define the exact relative forward errors for these solutions as $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$. Table 2 presents the calculated exact relative forward errors and the corresponding upper bound $U_{CL}$ (calculated using Equation 23) for these solutions across different values of $t$ and $\epsilon$.
$t$ $\epsilon$ $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$ $U_{CL}$
$1$ $7.0103\times 10^{-13}$ $1.3055\times 10^{-12}$ $2.7846\times 10^{-11}$
$7.0580\times 10^{-10}$ $1.3349\times 10^{-9}$ $2.8036\times 10^{-8}$
$8.9885\times 10^{-7}$ $1.81513\times 10^{-6}$ $3.5704\times 10^{-6}$
$5$ $7.1245\times 10^{-11}$ $1.1503\times 10^{-10}$ $9.5002\times 10^{-9}$
$8.9399\times 10^{-9}$ $2.0884\times 10^{-8}$ $1.1921\times 10^{-6}$
$1.0173\times 10^{-8}$ $1.5357\times 10^{-8}$ $1.3565\times 10^{-6}$
$9$ $7.1497\times 10^{-12}$ $1.2430\times 10^{-11}$ $1.3370\times 10^{-9}$
$6.6856\times 10^{-10}$ $1.2308\times 10^{-9}$ $1.2502\times 10^{-7}$
$1.0212\times 10^{-7}$ $2.5587\times 10^{-7}$ $1.9096\times 10^{-5}$
Table 2: Comparison of relative forward errors and their upper bounds for the complex solution of a perturbed RBLSE problem (11)
Table 2 shows that the exact relative forward errors of the complex solution to the RBLSE problem (11) are consistently lower than their respective upper bounds across various values of $t$ and $\epsilon$. This verifies the reliability of the derived upper bound $U_{CL}$ for the relative forward error.
Example 5.3.
Let $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}$. Let
$A_{i}$ $=$ $randn(m,n)\in{\mathbb{R}}^{m\times n},\;\textrm{for}\;i=0:3,$
$B_{i}$ $=$ $randn(m,d)\in{\mathbb{R}}^{m\times d},\;\textrm{for}\;i=0:3,$
$C_{i}$ $=$ $randn(p,n)\in{\mathbb{R}}^{p\times n},\;\textrm{for}\;i=0:3,$
$D_{i}$ $=$ $randn(p,d)\in{\mathbb{R}}^{p\times d},\;\textrm{for}\;i=0:3.$
Take $m=30t$, $n=10t$, $p=2t$, and $d=2$. Here, $t$ is an arbitrary number. We apply Algorithm 2 to determine the real solution of the RBLSE problem. Let $X_{RL}$ be the real solution of the RBLSE problem
$\min_{X,R_{r}}\left\lVert R_{r}\right\rVert_{F}\;\;\textrm{subject to}\;\;AX=B+R_{r},\;CX=D.$
Let $\epsilon_{1}=\left\lVert AX_{RL}-(B+R_{r})\right\rVert_{F}$ and $\epsilon_{2}=\left\lVert CX_{RL}-D\right\rVert_{F}$. In Table 3, we compute $\epsilon_{1}$ and $\epsilon_{2}$ for different values of $t$.
$t$ $\epsilon_{1}$ $\epsilon_{2}$
$1$ $2.9063\times 10^{-15}$ $3.0851\times 10^{-15}$
$3$ $2.5053\times 10^{-15}$ $5.5184\times 10^{-15}$
$5$ $3.7915\times 10^{-15}$ $1.0949\times 10^{-14}$
$7$ $3.9204\times 10^{-15}$ $1.3185\times 10^{-14}$
$9$ $5.2026\times 10^{-15}$ $1.7247\times 10^{-14}$
Table 3: Computational accuracy of Algorithm 2 for computing the real solution of the RBLSE problem (29)
Table 3 shows that the errors $\epsilon_{1}$ and $\epsilon_{2}$ across different values of $t$ are consistently below $10^{-14}$. This indicates that Algorithm 2 is highly effective in determining the real solution for the RBLSE problem (29).
Example 5.4.
In Example 5.3, we introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively, to examine how these small perturbations affect the real solution $X_{RL}$ of the RBLSE problem (29). These perturbations are measured normwise by the smallest $\epsilon$, as in (38). We define the exact relative forward errors for these solutions as $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$. Table 4 presents the computed exact relative forward errors and the corresponding upper bound $U_{RL}$ (calculated using Equation 40) for these solutions across different values of $t$ and $\epsilon$.
$t$ $\epsilon$ $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$ $U_{RL}$
$1$ $7.3638\times 10^{-12}$ $2.3814\times 10^{-11}$ $8.0339\times 10^{-10}$
$7.3677\times 10^{-10}$ $2.1476\times 10^{-9}$ $8.0381\times 10^{-8}$
$7.9614\times 10^{-8}$ $7.6845\times 10^{-7}$ $8.6858\times 10^{-6}$
$3$ $7.1742\times 10^{-14}$ $2.6216\times 10^{-13}$ $9.0524\times 10^{-12}$
$8.1702\times 10^{-11}$ $1.5855\times 10^{-10}$ $1.0309\times 10^{-8}$
$7.1758\times 10^{-9}$ $2.6438\times 10^{-8}$ $9.0544\times 10^{-7}$
$5$ $7.2567\times 10^{-11}$ $2.4717\times 10^{-10}$ $1.0430\times 10^{-8}$
$7.2735\times 10^{-9}$ $2.5130\times 10^{-8}$ $1.0454\times 10^{-6}$
$8.7538\times 10^{-7}$ $1.8374\times 10^{-6}$ $1.2581\times 10^{-4}$
Table 4: Comparison of relative forward errors and their upper bounds for the real solution of a perturbed RBLSE problem (29)
Table 4 shows that the exact relative forward errors of the real solution to the RBLSE problem (29) are consistently lower than their respective upper bounds across various values of $t$ and $\epsilon$. This verifies the reliability of the derived upper bound $U_{RL}$ for the relative forward error.
Example 5.5.
Let $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times 10}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{3\times 10}$, and $X_{0}=rand(10,3)+rand(10,3)\textit{{i}}\in{\mathbb{C}}^{10\times 3}$. Let
$M_{i}$ $=$ $rand(m,10)+rand(m,10)\textit{{i}}\in{\mathbb{C}}^{m\times 10},\;\textrm{for}\;i=1,2$
$R_{i}$ $=$ $rand(3,10)+rand(3,10)\textit{{i}}\in{\mathbb{C}}^{3\times 10},\;\textrm{for}\;i=1,2.$
Take $B=AX_{0}$ and $D=CX_{0}$. Thus, $X_{0}$ is the complex solution to the RBLSE problem (11). To measure the performance of the proposed technique for the RBLSE problem (11), Algorithm 1 is employed to determine the complex solution $X$. Let the error be $\epsilon=\left\lVert X-X_{0}\right\rVert_{F}$. Here, $m$ is a variable parameter. We evaluate the errors $\epsilon$ for various values of $m$. The relationship between the errors $\epsilon$ and $m$ is presented in Table 5.
$m$ $\epsilon=\|X-X_{0}\|_{F}$
$100$ $1.3154\times 10^{-14}$
$200$ $6.2150\times 10^{-15}$
$300$ $6.5603\times 10^{-15}$
$400$ $6.1485\times 10^{-15}$
$500$ $3.2441\times 10^{-15}$
Table 5: Computational accuracy of Algorithm 1 for computing the complex solution of the RBLSE problem (11)
Table 5 shows that the error $\epsilon$ between the complex solution derived from Algorithm 1 and the corresponding true solution to the RBLSE problem (11) remains consistently below $10^{-14}$ for various values of $m$. This indicates the high accuracy of Algorithm 1 in computing the complex solution to the RBLSE problem (11).
Example 5.6.
Let $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times 50}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{10\times 50}$, and $X_{0}=randn(50,30)\in{\mathbb{R}}^{50\times 30}$. Let
$A_{i}$ $=$ $randn(m,50)\in{\mathbb{R}}^{m\times 50},\;\textrm{for}\;i=0:3$
$C_{i}$ $=$ $randn(10,50)\in{\mathbb{R}}^{10\times 50},\;\textrm{for}\;i=0:3.$
Take $B=AX_{0}$ and $D=CX_{0}$. Clearly, $X_{0}$ is the real solution to the RBLSE problem (29). To measure the performance of the proposed technique for the RBLSE problem (29), Algorithm 2 is employed to determine the real solution $X$. Define the error as $\epsilon=\left\lVert X-X_{0}\right\rVert_{F}$. Here, $m$ is a variable parameter. We evaluate the errors $\epsilon$ for various values of $m$. The relationship between the errors $\epsilon$ and $m$ is presented in Table 6.
$m$ $\epsilon=\|X-X_{0}\|_{F}$
$1000$ $3.8948\times 10^{-14}$
$2000$ $4.4732\times 10^{-14}$
$3000$ $4.1257\times 10^{-14}$
$4000$ $3.7531\times 10^{-14}$
$5000$ $4.6532\times 10^{-14}$
Table 6: Computational accuracies of Algorithm 2 for computing the real solution of the RBLSE problem (29)
Table 6 shows that the error $\epsilon$ between the real solution derived from Algorithm 2 and the corresponding true solution to the RBLSE problem (29) remains consistently below $10^{-14}$ for various values of $m$. This indicates the high accuracy of Algorithm 2 in computing the real solution to the RBLSE problem (29).
6 Conclusions
We have presented an algebraic method for solving the RBLSE problem by transforming it into complex and real LSE problems. This transformation leverages the complex and real representations of reduced biquaternion matrices, enabling efficient computation of both complex and real solutions to the RBLSE problem. Additionally, we have calculated the upper bound for the relative forward error associated with these solutions, demonstrating the robustness and accuracy of our method in addressing the RBLSE problem.
7 Declarations
Author’s Contributions
The first author played a crucial role in overseeing the project and providing valuable guidance. Neha Bhadala made significant contributions at various stages of the research, including conceptualization, methodology design, software implementation, validation, and formal analysis. Additionally, Neha Bhadala took the lead in drafting the original manuscript and actively participated in the subsequent writing and editing phases.
Acknowledgement
Neha Bhadala expresses gratitude to the Government of India for funding her research under the Prime Minister’s Research Fellowship program (PMRF ID: $210059$).

### v2

@@
-On solutions of reduced biquaternion equality constrained least squares problem and their relative forward error bound 22footnotemark: 2
+On the specific solutions of reduced biquaternion equality constrained least squares problem and their relative forward error bound 22footnotemark: 2
 Abstract
 This study focuses on addressing the challenge of solving the reduced biquaternion equality constrained least squares (RBLSE) problem.
-We develop algebraic techniques to derive both complex and real solutions for the RBLSE problem by utilizing the complex and real forms of reduced biquaternion matrices.
-Additionally, we conduct a perturbation analysis for the RBLSE problem and establish an upper bound for the relative forward error of these solutions.
-Numerical examples are presented to illustrate the effectiveness of the proposed approaches and to verify the accuracy of the established upper bound for the relative forward errors.
+We develop algebraic techniques to derive real and complex solutions for the RBLSE problem by utilizing the real and complex forms of reduced biquaternion matrices.
+Furthermore, we propose algorithms and provide a detailed analysis of their computational complexity for finding special solutions to the RBLSE problem.
+A perturbation analysis is conducted, establishing an upper bound for the relative forward error of these solutions.
+This analysis ensures the accuracy and stability of the solutions in the presence of data perturbations, which is crucial for practical applications where errors arising from input inaccuracies can cause deviations between computed and true solutions.
+Numerical examples are presented to validate the proposed algorithms, demonstrate their effectiveness, and verify the accuracy of the established upper bound for the relative forward errors.
+These findings lay the groundwork for exploring applications in 3D and 4D algebra such as robotics, signal, and image processing, expanding their impact on practical and emerging domains.
 Keywords. Real representation matrix, Reduced biquaternion, Complex representation matrix, Equality constrained least squares problem.
@@
 1 Introduction
-Reduced biquaternions, first introduced by Segre in $1892$, have become fundamental in various domains such as signal processing, image analysis, and neural networks [4, 5, 6, 7].
+Reduced biquaternions, first introduced by Segre in $1892$, have become fundamental in various domains such as signal processing, image processing, and neural networks [5, 6, 7, 9].
 Unlike standard quaternions, reduced biquaternions, often referred to as commutative quaternions, possess the unique property of commutative multiplication.
 This property not only simplifies mathematical operations but also significantly boosts computational efficiency, as demonstrated in [12, 13].
+Reduced biquaternions also exhibit significant differences in their algebraic structure compared to traditional quaternions. Two notable elements, $e_{1}$ and $e_{2}$, are defined as:
+$e_{1}=\frac{1+\textit{{j}}}{2},\quad e_{2}=\frac{1-\textit{{j}}}{2}.$
+These elements satisfy $e_{1}e_{2}=0$, $e_{1}^{2}=e_{1}$, $e_{2}^{2}=e_{2}$, and are idempotent as well as divisors of zero.
+Any reduced biquaternion element of the form $c_{1}e_{1}$ or $c_{2}e_{2}$ (where $c_{1}$ and $c_{2}$ are complex numbers) also acts as a divisor of zero, highlighting the non-division algebra property of reduced biquaternions.
+Consequently, $\mathbb{Q}_{\mathbb{R}}$, the reduced biquaternion algebra, does not form a complete division algebra.
+The concept of units in $\mathbb{Q}_{\mathbb{R}}$ is also unique.
+An element $r\in\mathbb{Q}_{\mathbb{R}}$ is defined as a unit if there exists another element $s\in\mathbb{Q}_{\mathbb{R}}$ such that $rs=sr=1$.
+However, not every nonzero element in $\mathbb{Q}_{\mathbb{R}}$ is a unit, unlike in ${\mathbb{R}}$, ${\mathbb{C}}$ or in quaternions.
+For instance, $e_{1}$ or $e_{2}$ is not a unit in $\mathbb{Q}_{\mathbb{R}}$.
+These differences underline the distinct algebraic framework of reduced biquaternions and motivate their applicability to specific problems in mathematical and physical sciences.
+While reduced biquaternions have several advantages, their non-division algebra nature has negligible impact on practical applications.
+For instance, in [6, 9, 12, 13], the authors demonstrated the superior performance of reduced biquaternions over traditional quaternions in signal and image processing.
+This makes reduced biquaternions particularly advantageous for real-time applications in these domains.
+Moreover, their applicability extends beyond numerical computations to areas such as theoretical physics, including their connection to Maxwell’s equations [1, 10].
 The wide-ranging applications of reduced biquaternions have spurred significant research into both the theoretical and numerical aspects of matrices and matrix equations involving commutative quaternions.
@@
 The equality constrained least squares problem is a further study of the least squares problem, and in [17], the authors discussed solution techniques for computing the reduced biquaternion solution of the reduced biquaternion equality constrained least squares (RBLSE) problem.
-In this manuscript, we advance the research by developing methods to obtain both complex and real solutions for the RBLSE problem.
-Additionally, we establish an upper bound for the relative forward error associated with these solutions, ensuring the accuracy and reliability of our method in addressing RBLSE problems.
+Recent work has extended these studies to quaternions and split quaternions, proposing algorithms for solving split quaternion and quaternion equality constrained least squares problems [14, 20].
+Despite these advancements, existing studies primarily focus on computational techniques for finding general solutions to LSE problems over quaternions, split quaternions, and reduced biquaternions.
+However, specialized solutions to the RBLSE problem—such as real and complex solutions—have not been explored.
+Furthermore, the perturbation analysis of these solutions have not been addressed.
+In numerical analysis, the concept of relative forward error is central to evaluating the accuracy and stability of computed solutions, particularly for the RBLSE problem.
+This metric is crucial in the presence of data perturbations, where errors from machine precision, floating-point arithmetic, or input inaccuracies can lead to deviations between computed and true solutions.
+The relative forward error quantifies these deviations, providing insights into the sensitivity of solutions to data perturbations and enabling the assessment of solution accuracy and reliability.
+Understanding these errors is essential in practical applications, such as robotics, image processing, and control systems, where reduced biquaternion algebra plays a key role.
+To address these gaps, this study aims to:
+  1. 1.
+    Develop algebraic techniques to obtain real and complex solutions of the RBLSE problem.
+  2. 2.
+    Establish upper bounds for the relative forward error associated with these solutions to ensure accuracy and stability.
+  3. 3.
+    Propose efficient algorithms for solving the RBLSE problem and analyze their computational complexity.
+  4. 4.
+    Conduct numerical experiments to validate the proposed methods, assess their computational efficiency, and verify the accuracy of the established upper bounds.
+Among the proposed solutions, the computational complexity analysis reveals that finding the real solution to the RBLSE problem requires less computational time than obtaining the complex solution.
+Numerical experiments validate the computational efficiency of the algorithms, while the established upper bounds for the relative forward error provide insights into solution sensitivity and reliability under perturbations.
+These contributions are vital for advancing the numerical analysis of reduced biquaternions and ensuring the robustness of solutions in applications such as signal processing, color image processing, and robotics.
+By addressing both computational techniques and perturbation analysis, this study lays the groundwork for further exploration of reduced biquaternion algebra in practical and emerging domains.
 The organization of the manuscript is as follows: Section 2 covers the notation and preliminary concepts.
-Section 3 outlines the method for finding complex solution of the RBLSE problem, while Section 4 focuses on the technique for finding real solutions to the same problem.
+Section 3 outlines the method for finding the real solution of the RBLSE problem, while Section 4 focuses on the technique for finding complex solutions to the same problem.
 Lastly, Section 5 presents the numerical validation of the proposed methods.
@@
 In this paper, we adopt the following notations: $\mathbb{Q}_{\mathbb{R}}^{m\times n}$, ${\mathbb{C}}^{m\times n}$, and ${\mathbb{R}}^{m\times n}$ represent the sets of all $m\times n$ reduced biquaternion, complex, and real matrices.
-The notations $\left\lVert\cdot\right\rVert_{2}$ and $\left\lVert\cdot\right\rVert_{F}$ refer to the $2$-norm and Frobenius norm, respectively.
+The notation $\left\lVert\cdot\right\rVert_{F}$ represents the Frobenius norm, which applies to matrices over real numbers, complex numbers, or reduced biquaternions, while $\left\lVert\cdot\right\rVert_{2}$ denotes the $2$-norm, specifically for real or complex matrices.
 For matrices $A\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$ and $B\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $\left[A,B\right]$ represents their concatenation, i.e., $\begin{bmatrix}A&B\end{bmatrix}\in\mathbb{Q}_{\mathbb{R}}^{m\times(n+d)}$.
@@
 Additionally, $A^{-1}$ indicates the inverse of $A$ if it exists.
-The Matlab functions $randn(m,n)$ and $rand(m,n)$ produce $m\times n$ matrices filled with normally distributed and uniformly distributed random values, respectively.
+The Matlab function $rand(m,n)$ generates an $m\times n$ matrix filled with uniformly distributed random values.
 The abbreviations RBLSE (reduced biquaternion equality constrained least squares) and LSE (equality constrained least squares) will be used throughout this paper.
@@
 The real and complex representations of matrix $M$, denoted as $M^{R}$ and $M^{C}$, are given by:
-$M^{R}=\begin{bmatrix}M_{0}&-M_{1}&M_{2}&-M_{3}\\ M_{1}&M_{0}&M_{3}&M_{2}\\ M_{2}&-M_{3}&M_{0}&-M_{1}\\ M_{3}&M_{2}&M_{1}&M_{0}\end{bmatrix},\;\;M^{C}=\begin{bmatrix}N_{1}&N_{2}\\ N_{2}&N_{1}\end{bmatrix}.$ (2)
+$M^{R}=\begin{bmatrix}M_{0}&-M_{1}&M_{2}&-M_{3}\\ M_{1}&M_{0}&M_{3}&M_{2}\\ M_{2}&-M_{3}&M_{0}&-M_{1}\\ M_{3}&M_{2}&M_{1}&M_{0}\end{bmatrix},\quad M^{C}=\begin{bmatrix}N_{1}&N_{2}\\ N_{2}&N_{1}\end{bmatrix}.$ (2)
 Let $M_{c}^{R}$ represent the first block column of the matrix $M^{R}$.
@@
 where
-$Q_{m}=\begin{bmatrix}0&-I_{m}&0&0\\ I_{m}&0&0&0\\ 0&0&0&-I_{m}\\ 0&0&I_{m}&0\end{bmatrix},\;R_{m}=\begin{bmatrix}0&0&I_{m}&0\\ 0&0&0&I_{m}\\ I_{m}&0&0&0\\ 0&I_{m}&0&0\end{bmatrix},\mbox{and}\;S_{m}=\begin{bmatrix}0&0&0&-I_{m}\\ 0&0&I_{m}&0\\ 0&-I_{m}&0&0\\ I_{m}&0&0&0\end{bmatrix}.$ (4)
-We have
+$Q_{m}=\begin{bmatrix}0&-I_{m}&0&0\\ I_{m}&0&0&0\\ 0&0&0&-I_{m}\\ 0&0&I_{m}&0\end{bmatrix},\quad R_{m}=\begin{bmatrix}0&0&I_{m}&0\\ 0&0&0&I_{m}\\ I_{m}&0&0&0\\ 0&I_{m}&0&0\end{bmatrix},\quad S_{m}=\begin{bmatrix}0&0&0&-I_{m}\\ 0&0&I_{m}&0\\ 0&-I_{m}&0&0\\ I_{m}&0&0&0\end{bmatrix}.$ (4)
+The Frobenius norm, as described in [16, 17], is defined using the real representation of reduced biquaternion matrices.
+This definition aligns with the definition provided in (1), ensuring consistency.
+Specifically, the Frobenius norm of $M$ satisfies:
 $\left\lVert M\right\rVert_{F}=\frac{1}{2}\left\lVert M^{R}\right\rVert_{F}=\left\lVert M^{R}_{c}\right\rVert_{F}.$ (5)
@@
 Let $M_{c}^{C}$ represent the first block column of the matrix $M^{C}$.
-More precisely, $M_{c}^{C}$ is defined as: $M^{C}_{c}=\left[M_{1}^{T},M_{2}^{T}\right]^{T}.$ The matrix $M^{C}$ can be represented using $M^{C}_{c}$ as:
+More precisely, $M_{c}^{C}$ is defined as: $M^{C}_{c}=\left[N_{1}^{T},N_{2}^{T}\right]^{T}.$ The matrix $M^{C}$ can be represented using $M^{C}_{c}$ as:
 $M^{C}=\left[M_{c}^{C},P_{m}M_{c}^{C}\right],$ (6)
@@
 $P_{m}=\begin{bmatrix}0&I_{m}\\ I_{m}&0\end{bmatrix}.$ (7)
-We have
+The Frobenius norm, as described in [16, 17], is also expressed using the complex representation of reduced biquaternion matrices.
+This definition is consistent with the previously stated definition (1).
+Specifically, the Frobenius norm of $M$ satisfies:
 $\left\lVert M\right\rVert_{F}=\frac{1}{\sqrt{2}}\left\lVert M^{C}\right\rVert_{F}=\left\lVert M^{C}_{c}\right\rVert_{F}.$ (8)
@@
     $\left(PR\right)^{R}=P^{R}R^{R}$, $\left(PR\right)^{C}=P^{C}R^{C}$.
-3 An algebraic technique for complex solution of RBLSE problem
-This section focuses on an algebraic approach to derive the complex solution for the RBLSE problem. The method is based on analyzing the solution of the associated complex LSE problem. Suppose
-$A$ $=$ $M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\;\;\;\;B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (9)
-$C$ $=$ $R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\;\;\;\;D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (10)
-We will limit our discussion to the scenario where $m\geq n+d$, and both matrices $C$ and $C^{C}_{c}$ have full row rank. The RBLSE problem can then be stated as follows:
-$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;CX=D.$ (11)
-Consider a complex LSE problem
-$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{C}_{c}X=D^{C}_{c}.$ (12)
-Let the QR factorization of $\left(C^{C}_{c}\right)^{H}$ be as follows:
-$\left(C^{C}_{c}\right)^{H}=\widetilde{Q}\begin{bmatrix}\widetilde{R}\\ 0\end{bmatrix},$ (13)
-where $\widetilde{Q}\in{\mathbb{C}}^{n\times n}$ is a unitary matrix and $\widetilde{R}\in{\mathbb{C}}^{2p\times 2p}$ is a nonsingular upper triangular matrix.
-Now, partition $A^{C}_{c}\widetilde{Q}$ as:
-$A^{C}_{c}\widetilde{Q}=\left[\widetilde{P}_{1},\widetilde{P}_{2}\right],$ (14)
-where $\widetilde{P}_{1}\in{\mathbb{C}}^{2m\times 2p}$ and $\widetilde{P}_{2}\in{\mathbb{C}}^{2m\times(n-2p)}$.
-Theorem 3.1.
-Consider the RBLSE problem outlined in (11) and the complex LSE problem in (12), with notations specified in (13) and (14).
-For a matrix $X\in{\mathbb{C}}^{n\times d}$, $X$ is a complex solution of the RBLSE problem (11) if and only if $X$ solves the complex LSE problem (12).
-In this scenario, the unique solution $X$ with minimum norm can be expressed as:
-$X=\widetilde{Q}\begin{bmatrix}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}\\ \widetilde{P}_{2}^{\dagger}\left(B^{C}_{c}-\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}\right)\end{bmatrix}.$ (15)
+In the context of solving constrained least squares problems, the following result provides the closed-form solution for the complex LSE problem.
+This theorem, based on the QR factorization technique, is foundational for deriving efficient algorithms for more generalized problems discussed later in this paper.
+The result is formally stated as follows:
+Lemma 2.4.
+[8] Consider the complex LSE problem:
+$\min_{X}\|AX-B\|_{F}\quad\text{subject to}\quad CX=D,$ (9)
+where $A\in\mathbb{C}^{m\times n}$, $B\in\mathbb{C}^{m\times d}$, $C\in\mathbb{C}^{p\times n}$, and $D\in\mathbb{C}^{p\times d}$.
+Assume that the matrix $C$ has full row rank.
+Let the QR factorization of $C^{H}$ be given by:
+$C^{H}=Q\begin{bmatrix}R\\ 0\end{bmatrix},$ (10)
+where $Q\in\mathbb{C}^{n\times n}$ is a unitary matrix and $R\in\mathbb{C}^{p\times p}$ is a nonsingular upper triangular matrix. Further, partition the product $AQ$ as:
+$AQ=\left[P_{1},P_{2}\right],$ (11)
+where $P_{1}\in\mathbb{C}^{m\times p}$ and $P_{2}\in\mathbb{C}^{m\times(n-p)}$.
+The unique solution $X\in\mathbb{C}^{n\times d}$ with minimum norm for the complex LSE problem (9) is given by:
+$X=Q\begin{bmatrix}\left(R^{H}\right)^{-1}D\\ P_{2}^{\dagger}\left(B-P_{1}\left(R^{H}\right)^{-1}D\right)\end{bmatrix}.$ (12)
 Proof.
-If $X\in{\mathbb{C}}^{n\times d}$ is a solution of the complex LSE problem (12), then
-$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\textrm{min},\;\;C^{C}_{c}X=D^{C}_{c}.$ (16)
-The Frobenius norm of a complex matrix remains invariant under unitary transformations. Since the matrix $P_{m}$ in (7) is unitary, it follows that:
-$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\left\lVert P_{m}\left(A^{C}_{c}X-B^{C}_{c}\right)\right\rVert_{F}.$
-Using equations (2), (6), (8), along with Lemmas 2.2 and 2.3, we obtain
-$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}$ $=$ $\frac{1}{\sqrt{2}}\left\lVert\left[\left(A^{C}_{c}X-B^{C}_{c}\right),P_{m}\left(A^{C}_{c}X-B^{C}_{c}\right)\right]\right\rVert_{F}$
-$=$ $\frac{1}{\sqrt{2}}\left\lVert\left[A^{C}_{c}X,P_{m}A^{C}_{c}X\right]-\left[B^{C}_{c},P_{m}B^{C}_{c}\right]\right\rVert_{F}$
-$=$ $\frac{1}{\sqrt{2}}\left\lVert\left[A^{C}_{c},P_{m}A^{C}_{c}\right]\begin{bmatrix}X&0\\ 0&X\end{bmatrix}-\left[B^{C}_{c},P_{m}B^{C}_{c}\right]\right\rVert_{F}$
-$=$ $\frac{1}{\sqrt{2}}\left\lVert A^{C}X^{C}-B^{C}\right\rVert_{F}$
-$=$ $\frac{1}{\sqrt{2}}\left\lVert\left(AX-B\right)^{C}\right\rVert_{F}$
-$=$ $\left\lVert AX-B\right\rVert_{F}.$
-From (16), we obtain
-$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\left\lVert AX-B\right\rVert_{F}=\textrm{min},$ (17)
-and
-$\left[C^{C}_{c},P_{p}C^{C}_{c}\right]\begin{bmatrix}X&0\\ 0&X\end{bmatrix}=\left[D^{C}_{c},P_{p}D^{C}_{c}\right].$
-Using (6), we know that $C^{C}=\left[C^{C}_{c},P_{p}C^{C}_{c}\right]$ and $D^{C}=\left[D^{C}_{c},P_{p}D^{C}_{c}\right]$. Applying this, we get
-$C^{C}X^{C}$ $=$ $D^{C},$
-$\left(CX\right)^{C}$ $=$ $D^{C},$
-$CX$ $=$ $D.$ (18)
-By combining (17) and (18), we conclude that $X\in{\mathbb{C}}^{n\times d}$ is a complex solution to the RBLSE problem (11), and vice versa.
-To find the expression for $X$, we solve the complex LSE problem (12). Set
-$\widetilde{Q}^{H}X=\begin{bmatrix}Y\\ Z\end{bmatrix},$
-where $Y\in{\mathbb{C}}^{2p\times d}$ and $Z\in{\mathbb{C}}^{(n-2p)\times d}$. Equation 12 can be rewritten as
-$\min_{X}\left\lVert A^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X-B^{C}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X=D^{C}_{c}.$ (19)
-Utilizing (13), we have
-$C^{C}_{c}\widetilde{Q}\widetilde{Q}^{H}X=D^{C}_{c}\implies\begin{bmatrix}\widetilde{R}^{H}&0\end{bmatrix}\begin{bmatrix}Y\\ Z\end{bmatrix}=D^{C}_{c}.$
-Since $\widetilde{R}^{H}$ is a nonsingular matrix, we get $Y=\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$. Using (14), Equation 19 takes the form
-$\min_{Z}\left\lVert\widetilde{P}_{2}Z-\left(B^{C}_{c}-\widetilde{P}_{1}Y\right)\right\rVert_{F}.$
-The minimum norm solution of the above least squares problem is $Z=\widetilde{P}_{2}^{\dagger}\left(B^{C}_{c}-\widetilde{P}_{1}Y\right)$. Thus, we can derive the desired expression for $X$. ∎
-Next, we aim to examine how perturbations in $A$, $B$, $C$, and $D$ affect the complex solution $X_{CL}$ of the RBLSE problem (11). Let
-$\widehat{A}=A+\Delta A,\;\widehat{B}=B+\Delta B,\;\widehat{C}=C+\Delta C,\;\textrm{and}\;\widehat{D}=D+\Delta D,$ (20)
-where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively.
-We assume that the perturbation $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to guarantee that the perturbed matrices $\widehat{C}$ and $\widehat{C}^{C}_{c}$ retain full row rank.
-These perturbations are measured normwise by the smallest $\epsilon$ for which
-$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\left\lVert\Delta C\right\rVert_{F}\leq\epsilon\left\lVert C\right\rVert_{F},\;\textrm{and}\;\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (21)
-Let $\widehat{X}_{CL}$ be the complex solution to the perturbed RBLSE problem
-$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\;\;\textrm{subject to}\;\;\widehat{C}X=\widehat{D},$ (22)
-and let $\Delta X_{CL}=\widehat{X}_{CL}-X_{CL}$.
-Theorem 3.2.
-Consider the RBLSE problem outlined in (11) and the perturbed RBLSE problem described in (22).
-If the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are sufficiently small, as described in (21), then we have
-$\begin{split}\frac{\left\lVert\Delta X_{CL}\right\rVert_{F}}{\left\lVert X_{CL}\right\rVert_{F}}\leq\epsilon&\left(\mathcal{K}_{A}^{C}\left(\frac{\left\lVert D^{C}_{c}\right\rVert_{F}}{\left\lVert C^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}+1\right)+\mathcal{K}_{B}^{C}\left(\frac{\left\lVert B^{C}_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}+1\right)\right.\\ &\left.+\left(\mathcal{K}_{B}^{C}\right)^{2}\left(\frac{\left\lVert C^{C}_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}}\left\lVert A^{C}_{c}\mathcal{L}_{c}\right\rVert_{2}+1\right)\frac{\left\lVert R_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}\right)+O(\epsilon^{2})\equiv U_{CL},\end{split}$ (23)
-where
-$\mathcal{K}_{B}^{C}$ $=$ $\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert\left(A^{C}_{c}P_{c}\right)^{\dagger}\right\rVert_{2},\;\;\;\mathcal{K}_{A}^{C}=\left\lVert C^{C}_{c}\right\rVert_{F}\left\lVert\mathcal{L}_{c}\right\rVert_{2},\;\;\;\mathcal{L}_{c}=\left(I_{n}-\left(A^{C}_{c}P_{c}\right)^{\dagger}A^{C}_{c}\right)\left(C^{C}_{c}\right)^{\dagger},\;\;\;\;$
-$P_{c}$ $=$ $I_{n}-\left(C^{C}_{c}\right)^{\dagger}C^{C}_{c},\;\;\;\;\;\;\;\;\;\;\;R_{c}=B^{C}_{c}-A^{C}_{c}X_{CL}.$
-Proof.
-The perturbed complex LSE problem corresponding to the perturbed RBLSE problem (22) is given by:
-$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{C}_{c}X-\left(\widehat{B}\right)^{C}_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(\widehat{C}\right)^{C}_{c}X=\left(\widehat{D}\right)^{C}_{c}.$ (24)
-Using Theorem 3.1, we know that $\widehat{X}_{CL}$ is the solution to the perturbed complex LSE problem (24). From (20) and utilizing Lemma 2.3, we have
-$\left(\widehat{A}\right)^{C}_{c}=A^{C}_{c}+\left(\Delta A\right)^{C}_{c},\;\left(\widehat{B}\right)^{C}_{c}=B^{C}_{c}+\left(\Delta B\right)^{C}_{c},\;\left(\widehat{C}\right)^{C}_{c}=C^{C}_{c}+\left(\Delta C\right)^{C}_{c},\;\textrm{and}\;\left(\widehat{D}\right)^{C}_{c}=D^{C}_{c}+\left(\Delta D\right)^{C}_{c}.$
-Thus, the perturbed complex LSE problem (24) can be rewritten as:
-$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(A^{C}_{c}+\left(\Delta A\right)^{C}_{c}\right)X-\left(B^{C}_{c}+\left(\Delta B\right)^{C}_{c}\right)\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(C^{C}_{c}+\left(\Delta C\right)^{C}_{c}\right)X=\left(D^{C}_{c}+\left(\Delta D\right)^{C}_{c}\right).$ (25)
-Using (8) and (21), we can establish the following bounds for the perturbation:
-$\left\lVert\left(\Delta A\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{C}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta B\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{C}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta C\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{C}_{c}\right\rVert_{F},\;\textrm{and}\;\left\lVert\left(\Delta D\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{C}_{c}\right\rVert_{F}.$ (26)
-With the perturbed problem (25) and the bounds in (26), and using Theorem 3.1, the sensitivity analysis of the complex solution to the RBLSE problem (11) reduces to evaluating the sensitivity of the solution to the complex LSE problem (12).
-Consequently, the upper bound $U_{CL}$ for the relative forward error of the complex solution to the RBLSE problem can be obtained from [1, Equation $4.11$].
-∎
+Equation (9) can be rewritten as
+$\min_{X}\left\lVert AQQ^{H}X-B\right\rVert_{F}\quad\textrm{subject to}\quad CQQ^{H}X=D.$ (13)
+Set
+$Q^{H}X=\begin{bmatrix}Y\\ Z\end{bmatrix},$
+where $Y\in\mathbb{C}^{p\times d}$ and $Z\in\mathbb{C}^{(n-p)\times d}$. Utilizing (10), we have
+$CQQ^{H}X=D\implies\begin{bmatrix}R^{H}&0\end{bmatrix}\begin{bmatrix}Y\\ Z\end{bmatrix}=D.$
+Since $R^{H}$ is a nonsingular matrix, we get $Y=\left(R^{H}\right)^{-1}D$. Using (11), equation (13) takes the form
+$\min_{Z}\left\lVert P_{2}Z-\left(B-P_{1}Y\right)\right\rVert_{F}.$
+The minimum norm solution of the above least squares problem is $Z=P_{2}^{\dagger}\left(B-P_{1}Y\right)$.
+Finally, substituting $Y$ and $Z$ back into $Q^{H}X=\begin{bmatrix}Y\\ Z\end{bmatrix}$, we can derive the desired expression for $X$. This completes the proof. ∎
 3 An algebraic technique for real solution of RBLSE problem
 This section focuses on an algebraic approach to derive the real solution for the RBLSE problem. The method is based on analyzing the solution of the associated real LSE problem. Suppose
-$A$ $=$ $A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\;\;\;\;B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (27)
-$C$ $=$ $C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\;\;\;\;\;D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (28)
-We will limit our discussion to the scenario where $m\geq n+d$, and both matrices $C$ and $C^{R}_{c}$ have full row rank. The RBLSE problem can then be stated as follows:
-$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;CX=D.$ (29)
+$A$ $=$ $A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\quad B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (14)
+$C$ $=$ $C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\quad D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (15)
+We will limit our discussion to the scenario where $m\geq n+d$, and matrix $C^{R}_{c}$ has full row rank. The RBLSE problem can then be stated as follows:
+$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\quad\textrm{subject to}\quad CX=D.$ (16)
 Consider a real LSE problem
-$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{R}_{c}X=D^{R}_{c}.$ (30)
-Let the QR factorization of $\left(C^{R}_{c}\right)^{T}$ be as follows:
+$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}\quad\textrm{subject to}\quad C^{R}_{c}X=D^{R}_{c}.$ (17)
+Theorem 3.1.
+Consider the RBLSE problem outlined in (16) and the real LSE problem in (17). Let the QR factorization of $\left(C^{R}_{c}\right)^{T}$ be given by:
 $\left(C^{R}_{c}\right)^{T}=\bar{Q}\begin{bmatrix}\bar{R}\\ 0\end{bmatrix},$ (18)
-where $\bar{Q}\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix and $\bar{R}\in{\mathbb{R}}^{4p\times 4p}$ is a nonsingular upper triangular matrix. Now, partition $A^{R}_{c}\bar{Q}$ as:
+where $\bar{Q}\in\mathbb{R}^{n\times n}$ is an orthonormal matrix and $\bar{R}\in\mathbb{R}^{4p\times 4p}$ is a nonsingular upper triangular matrix.
+Further, partition the product $A^{R}_{c}\bar{Q}$ as:
 $A^{R}_{c}\bar{Q}=\left[\bar{P}_{1},\bar{P}_{2}\right],$ (19)
 where $\bar{P}_{1}\in\mathbb{R}^{4m\times 4p}$ and $\bar{P}_{2}\in\mathbb{R}^{4m\times(n-4p)}$.
-Theorem 4.1.
-Consider the RBLSE problem outlined in (29) and the real LSE problem in (30), with notations specified in (31) and (32).
-For a matrix $X\in{\mathbb{R}}^{n\times d}$, $X$ is a real solution of the RBLSE problem (29) if and only if $X$ solves the real LSE problem (30).
-In this scenario, the unique solution with minimum norm $X$ can be expressed as:
+Let $X\in\mathbb{R}^{n\times d}$.
+Then, $X$ is a real solution of the RBLSE problem (16) if and only if $X$ solves the real LSE problem (17).
+In this scenario, the unique solution $X$ with minimum norm can be expressed as:
 $X=\bar{Q}\begin{bmatrix}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}\\ \bar{P}_{2}^{\dagger}\left(B^{R}_{c}-\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}\right)\end{bmatrix}.$ (20)
@@
 If $X\in{\mathbb{R}}^{n\times d}$ is a solution of the real LSE problem (17), then
-$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\textrm{min},\;\;C^{R}_{c}X=D^{R}_{c}.$ (34)
+$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\textrm{min},\quad C^{R}_{c}X=D^{R}_{c}.$ (21)
 The Frobenius norm of a real matrix remains invariant under orthogonal transformations. Since the matrices $Q_{m}$, $R_{m}$, and $S_{m}$ in (4) are orthogonal, it follows that:
@@
 By combining (22) and (23), we conclude that $X\in{\mathbb{R}}^{n\times d}$ is a real solution to the RBLSE problem (16), and vice versa.
-The expression for $X$ can be obtained by following the proof method of Theorem 3.1.
-∎
+To determine the expression for $X$, we solve the real LSE problem given in (17). Using Lemma 2.4, we derive the explicit expression for $X$. ∎
 Next, we aim to examine how perturbations in $A$, $B$, $C$, and $D$ affect the real solution $X_{RL}$ of the RBLSE problem (16). Let
-$\widehat{A}=A+\Delta A,\;\widehat{B}=B+\Delta B,\;\widehat{C}=C+\Delta C,\;\textrm{and}\;\widehat{D}=D+\Delta D,$ (37)
+$\widehat{A}=A+\Delta A,\quad\widehat{B}=B+\Delta B,\quad\widehat{C}=C+\Delta C,\quad\textrm{and}\quad\widehat{D}=D+\Delta D,$ (24)
 where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively.
-We assume that the perturbation $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to ensure that the perturbed matrices $\widehat{C}$ and $\widehat{C}^{R}_{c}$ retain full row rank.
+We assume that the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to ensure that the perturbed matrix $\widehat{C}^{R}_{c}$ retains full row rank.
 These perturbations are measured normwise by the smallest $\epsilon$ for which
-$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\textrm{and}\;\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (38)
+$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\quad\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\quad\left\lVert\Delta C\right\rVert_{F}\leq\epsilon\left\lVert C\right\rVert_{F},\quad\textrm{and}\quad\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (25)
 Let $\widehat{X}_{RL}$ be the real solution to the perturbed RBLSE problem
-$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\;\;\textrm{subject to}\;\;\widehat{C}X=\widehat{D},$ (39)
+$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\quad\textrm{subject to}\quad\widehat{C}X=\widehat{D},$ (26)
 and let $\Delta X_{RL}=\widehat{X}_{RL}-X_{RL}$.
@@
 The perturbed real LSE problem corresponding to the perturbed RBLSE problem (26) is given by:
-$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{R}_{c}X-\left(\widehat{B}\right)^{R}_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(\widehat{C}\right)^{R}_{c}X=\left(\widehat{D}\right)^{R}_{c}.$ (41)
+$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{R}_{c}X-\left(\widehat{B}\right)^{R}_{c}\right\rVert_{F}\quad\textrm{subject to}\quad\left(\widehat{C}\right)^{R}_{c}X=\left(\widehat{D}\right)^{R}_{c}.$ (28)
 Using Theorem 3.1, we know that $\widehat{X}_{RL}$ is the solution to the perturbed real LSE problem (28). From (24) and utilizing Lemma 2.3, we have
-$\left(\widehat{A}\right)^{R}_{c}=A^{R}_{c}+\left(\Delta A\right)^{R}_{c},\;\left(\widehat{B}\right)^{R}_{c}=B^{R}_{c}+\left(\Delta B\right)^{R}_{c},\;\left(\widehat{C}\right)^{R}_{c}=C^{R}_{c}+\left(\Delta C\right)^{R}_{c},\;\textrm{and}\;\left(\widehat{D}\right)^{R}_{c}=D^{R}_{c}+\left(\Delta D\right)^{R}_{c}.$
+$\left(\widehat{A}\right)^{R}_{c}=A^{R}_{c}+\left(\Delta A\right)^{R}_{c},\quad\left(\widehat{B}\right)^{R}_{c}=B^{R}_{c}+\left(\Delta B\right)^{R}_{c},\quad\left(\widehat{C}\right)^{R}_{c}=C^{R}_{c}+\left(\Delta C\right)^{R}_{c},\quad\left(\widehat{D}\right)^{R}_{c}=D^{R}_{c}+\left(\Delta D\right)^{R}_{c}.$
 Thus, the perturbed real LSE problem (28) can be rewritten as:
-$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(A^{R}_{c}+\left(\Delta A\right)^{R}_{c}\right)X-\left(B^{R}_{c}+\left(\Delta B\right)^{R}_{c}\right)\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(C^{R}_{c}+\left(\Delta C\right)^{R}_{c}\right)X=\left(D^{R}_{c}+\left(\Delta D\right)^{R}_{c}\right).$ (42)
+$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(A^{R}_{c}+\left(\Delta A\right)^{R}_{c}\right)X-\left(B^{R}_{c}+\left(\Delta B\right)^{R}_{c}\right)\right\rVert_{F}\quad\textrm{subject to}\quad\left(C^{R}_{c}+\left(\Delta C\right)^{R}_{c}\right)X=\left(D^{R}_{c}+\left(\Delta D\right)^{R}_{c}\right).$ (29)
 Using (5) and (25), we can establish the following bounds for the perturbation:
-$\left\lVert\left(\Delta A\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{R}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta B\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{R}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta C\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{R}_{c}\right\rVert_{F},\;\textrm{and}\;\left\lVert\left(\Delta D\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{R}_{c}\right\rVert_{F}.$ (43)
+$\left\lVert\left(\Delta A\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{R}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta B\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{R}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta C\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{R}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta D\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{R}_{c}\right\rVert_{F}.$ (30)
 With the perturbed problem (29) and the bounds in (30), and using Theorem 3.1, the sensitivity analysis of the real solution to the RBLSE problem (16) reduces to evaluating the sensitivity of the solution to the real LSE problem (17).
-Consequently, the upper bound $U_{RL}$ for the relative forward error of the real solution to the RBLSE problem can be obtained from [1, Equation $4.11$].
+Consequently, the upper bound $U_{RL}$ for the relative forward error of the real solution to the RBLSE problem can be obtained from [2, Equation $4.11$].
+∎
+4 An algebraic technique for complex solution of RBLSE problem
+This section focuses on an algebraic approach to derive the complex solution for the RBLSE problem. The method is based on analyzing the solution of the associated complex LSE problem. Suppose
+$A$ $=$ $M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\quad B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (31)
+$C$ $=$ $R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\quad D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (32)
+We will limit our discussion to the scenario where $m\geq n+d$, and matrix $C^{C}_{c}$ has full row rank. The RBLSE problem can then be stated as follows:
+$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\quad\textrm{subject to}\quad CX=D.$ (33)
+Consider a complex LSE problem
+$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}\quad\textrm{subject to}\quad C^{C}_{c}X=D^{C}_{c}.$ (34)
+Theorem 4.1.
+Consider the RBLSE problem outlined in (33) and the complex LSE problem in (34). Let the QR factorization of $(C^{C}_{c})^{H}$ be given by:
+$\left(C^{C}_{c}\right)^{H}=\widetilde{Q}\begin{bmatrix}\widetilde{R}\\ 0\end{bmatrix},$ (35)
+where $\widetilde{Q}\in\mathbb{C}^{n\times n}$ is a unitary matrix and $\widetilde{R}\in\mathbb{C}^{2p\times 2p}$ is a nonsingular upper triangular matrix.
+Further, partition the product $A^{C}_{c}\widetilde{Q}$ as:
+$A^{C}_{c}\widetilde{Q}=\left[\widetilde{P}_{1},\widetilde{P}_{2}\right],$ (36)
+where $\widetilde{P}_{1}\in\mathbb{C}^{2m\times 2p}$ and $\widetilde{P}_{2}\in\mathbb{C}^{2m\times(n-2p)}$.
+Let $X\in\mathbb{C}^{n\times d}$.
+Then, $X$ is a complex solution of the RBLSE problem (33) if and only if $X$ solves the complex LSE problem (34).
+In this scenario, the unique solution $X$ with minimum norm can be expressed as:
+$X=\widetilde{Q}\begin{bmatrix}(\widetilde{R}^{H})^{-1}D^{C}_{c}\\ \widetilde{P}_{2}^{\dagger}\big{(}B^{C}_{c}-\widetilde{P}_{1}(\widetilde{R}^{H})^{-1}D^{C}_{c}\big{)}\end{bmatrix}.$ (37)
+Proof.
+If $X\in{\mathbb{C}}^{n\times d}$ is a solution of the complex LSE problem (34), then
+$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\textrm{min},\quad C^{C}_{c}X=D^{C}_{c}.$ (38)
+The Frobenius norm of a complex matrix remains invariant under unitary transformations. Since the matrix $P_{m}$ in (7) is unitary, it follows that:
+$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\left\lVert P_{m}\left(A^{C}_{c}X-B^{C}_{c}\right)\right\rVert_{F}.$
+Using equations (2), (6), (8), along with Lemmas 2.2 and 2.3, we obtain
+$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}$ $=$ $\frac{1}{\sqrt{2}}\left\lVert\left[\left(A^{C}_{c}X-B^{C}_{c}\right),P_{m}\left(A^{C}_{c}X-B^{C}_{c}\right)\right]\right\rVert_{F}$
+$=$ $\frac{1}{\sqrt{2}}\left\lVert\left[A^{C}_{c}X,P_{m}A^{C}_{c}X\right]-\left[B^{C}_{c},P_{m}B^{C}_{c}\right]\right\rVert_{F}$
+$=$ $\frac{1}{\sqrt{2}}\left\lVert\left[A^{C}_{c},P_{m}A^{C}_{c}\right]\begin{bmatrix}X&0\\ 0&X\end{bmatrix}-\left[B^{C}_{c},P_{m}B^{C}_{c}\right]\right\rVert_{F}$
+$=$ $\frac{1}{\sqrt{2}}\left\lVert A^{C}X^{C}-B^{C}\right\rVert_{F}$
+$=$ $\frac{1}{\sqrt{2}}\left\lVert\left(AX-B\right)^{C}\right\rVert_{F}$
+$=$ $\left\lVert AX-B\right\rVert_{F}.$
+From (38), we obtain
+$\left\lVert A^{C}_{c}X-B^{C}_{c}\right\rVert_{F}=\left\lVert AX-B\right\rVert_{F}=\textrm{min},$ (39)
+and
+$\left[C^{C}_{c},P_{p}C^{C}_{c}\right]\begin{bmatrix}X&0\\ 0&X\end{bmatrix}=\left[D^{C}_{c},P_{p}D^{C}_{c}\right].$
+Using (6), we know that $C^{C}=\left[C^{C}_{c},P_{p}C^{C}_{c}\right]$ and $D^{C}=\left[D^{C}_{c},P_{p}D^{C}_{c}\right]$. Applying this, we get
+$C^{C}X^{C}$ $=$ $D^{C},$
+$\left(CX\right)^{C}$ $=$ $D^{C},$
+$CX$ $=$ $D.$ (40)
+By combining (39) and (40), we conclude that $X\in{\mathbb{C}}^{n\times d}$ is a complex solution to the RBLSE problem (33), and vice versa.
+To find the expression for $X$, we solve the complex LSE problem (34). Using Lemma 2.4, we derive the explicit expression for $X$. ∎
+Next, we aim to examine how perturbations in $A$, $B$, $C$, and $D$ affect the complex solution $X_{CL}$ of the RBLSE problem (33). Let
+$\widehat{A}=A+\Delta A,\quad\widehat{B}=B+\Delta B,\quad\widehat{C}=C+\Delta C,\quad\textrm{and}\quad\widehat{D}=D+\Delta D,$ (41)
+where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively.
+We assume that the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to guarantee that the perturbed matrix $\widehat{C}^{C}_{c}$ retains full row rank.
+These perturbations are measured normwise by the smallest $\epsilon$ for which
+$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\quad\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\quad\left\lVert\Delta C\right\rVert_{F}\leq\epsilon\left\lVert C\right\rVert_{F},\quad\textrm{and}\quad\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (42)
+Let $\widehat{X}_{CL}$ be the complex solution to the perturbed RBLSE problem
+$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\quad\textrm{subject to}\quad\widehat{C}X=\widehat{D},$ (43)
+and let $\Delta X_{CL}=\widehat{X}_{CL}-X_{CL}$.
+Theorem 4.2.
+Consider the RBLSE problem outlined in (33) and the perturbed RBLSE problem described in (43).
+If the perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are sufficiently small, as described in (42), then we have
+$\begin{split}\frac{\left\lVert\Delta X_{CL}\right\rVert_{F}}{\left\lVert X_{CL}\right\rVert_{F}}\leq\epsilon&\left(\mathcal{K}_{A}^{C}\left(\frac{\left\lVert D^{C}_{c}\right\rVert_{F}}{\left\lVert C^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}+1\right)+\mathcal{K}_{B}^{C}\left(\frac{\left\lVert B^{C}_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}+1\right)\right.\\ &\left.+\left(\mathcal{K}_{B}^{C}\right)^{2}\left(\frac{\left\lVert C^{C}_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}}\left\lVert A^{C}_{c}\mathcal{L}_{c}\right\rVert_{2}+1\right)\frac{\left\lVert R_{c}\right\rVert_{F}}{\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert X_{CL}\right\rVert_{F}}\right)+O(\epsilon^{2})\equiv U_{CL},\end{split}$ (44)
+where
+$\mathcal{K}_{B}^{C}$ $=$ $\left\lVert A^{C}_{c}\right\rVert_{F}\left\lVert\left(A^{C}_{c}P_{c}\right)^{\dagger}\right\rVert_{2},\;\;\;\mathcal{K}_{A}^{C}=\left\lVert C^{C}_{c}\right\rVert_{F}\left\lVert\mathcal{L}_{c}\right\rVert_{2},\;\;\;\mathcal{L}_{c}=\left(I_{n}-\left(A^{C}_{c}P_{c}\right)^{\dagger}A^{C}_{c}\right)\left(C^{C}_{c}\right)^{\dagger},\;\;\;\;$
+$P_{c}$ $=$ $I_{n}-\left(C^{C}_{c}\right)^{\dagger}C^{C}_{c},\;\;\;\;\;\;\;\;\;\;\;R_{c}=B^{C}_{c}-A^{C}_{c}X_{CL}.$
+Proof.
+The perturbed complex LSE problem corresponding to the perturbed RBLSE problem (43) is given by:
+$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{C}_{c}X-\left(\widehat{B}\right)^{C}_{c}\right\rVert_{F}\quad\textrm{subject to}\quad\left(\widehat{C}\right)^{C}_{c}X=\left(\widehat{D}\right)^{C}_{c}.$ (45)
+Using Theorem 4.1, we know that $\widehat{X}_{CL}$ is the solution to the perturbed complex LSE problem (45). From (41) and utilizing Lemma 2.3, we have
+$\left(\widehat{A}\right)^{C}_{c}=A^{C}_{c}+\left(\Delta A\right)^{C}_{c},\quad\left(\widehat{B}\right)^{C}_{c}=B^{C}_{c}+\left(\Delta B\right)^{C}_{c},\quad\left(\widehat{C}\right)^{C}_{c}=C^{C}_{c}+\left(\Delta C\right)^{C}_{c},\quad\left(\widehat{D}\right)^{C}_{c}=D^{C}_{c}+\left(\Delta D\right)^{C}_{c}.$
+Thus, the perturbed complex LSE problem (45) can be rewritten as:
+$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(A^{C}_{c}+\left(\Delta A\right)^{C}_{c}\right)X-\left(B^{C}_{c}+\left(\Delta B\right)^{C}_{c}\right)\right\rVert_{F}\quad\textrm{subject to}\quad\left(C^{C}_{c}+\left(\Delta C\right)^{C}_{c}\right)X=\left(D^{C}_{c}+\left(\Delta D\right)^{C}_{c}\right).$ (46)
+Using (8) and (42), we can establish the following bounds for the perturbation:
+$\left\lVert\left(\Delta A\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{C}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta B\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{C}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta C\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{C}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta D\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{C}_{c}\right\rVert_{F}.$ (47)
+With the perturbed problem (46) and the bounds in (47), and using Theorem 4.1, the sensitivity analysis of the complex solution to the RBLSE problem (33) reduces to evaluating the sensitivity of the solution to the complex LSE problem (34).
+Consequently, the upper bound $U_{CL}$ for the relative forward error of the complex solution to the RBLSE problem can be obtained from [2, Equation $4.11$].
 ∎
 5 Numerical Verification
-Building on the previous discussions, this section presents numerical algorithms designed to find special solutions to the RBLSE problem.
-We also include numerical examples to validate these algorithms.
-Additionally, we assess the upper bound for the relative forward error of both complex and real solutions to the RBLSE problem.
+In this section, we present numerical algorithms for computing two special solutions (real and complex) of the RBLSE problem.
+Each algorithm is complemented with a detailed computational complexity analysis through a step-by-step flop count.
+Additionally, we validate these algorithms using numerical examples to assess their performance and reliability.
+Algorithm 1 For the real solution of the RBLSE problem (16)
+Input: $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that matrix $C^{R}_{c}$ has full row rank.
+Output: $X\in{\mathbb{R}}^{n\times d}$.
+  1. 1.
+    Find the QR factorization of $\left(C^{R}_{c}\right)^{T}$, as described in (18).
+  2. 2.
+    Partition the matrix $A^{R}_{c}\bar{Q}$ as shown in (19).
+  3. 3.
+    The real solution $X$ to the RBLSE problem (16) is given by (20).
 Algorithm 2 For the complex solution of the RBLSE problem (33)
-Input: $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that both matrices $C$ and $C^{C}_{c}$ have full row rank.
-Output: $X$.
+Input: $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that matrix $C^{C}_{c}$ has full row rank.
+Output: $X\in{\mathbb{C}}^{n\times d}$.
   1. 1.
@@
     The complex solution $X$ to the RBLSE problem (33) is given by (37).
-Algorithm 2 For the real solution of the RBLSE problem (29)
-Input: $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that both matrices $C$ and $C^{R}_{c}$ have full row rank.
-Output: $X$.
-  1. 1.
-    Find the QR factorization of $\left(C^{R}_{c}\right)^{T}$, as described in (31).
-  2. 2.
-    Partition the matrix $A^{R}_{c}\bar{Q}$ as shown in (32).
-  3. 3.
-    The real solution $X$ to the RBLSE problem (29) is given by (33).
+Next, to assess the performance of the algorithms, a step-by-step flop count is presented.
+This detailed analysis highlights the computational cost at each stage, providing insights into the efficiency of the algorithm.
+First, to analyze the efficiency of the real solution algorithm, a detailed step-by-step flop count is provided below.
+Step Description Flop Count
+1 QR decomposition of $\left(C^{R}_{c}\right)^{T}$ $O(32np^{2})$
+2 Compute $A^{R}_{c}\bar{Q}$ $O(8mn^{2}-4mn)$
+3 Partition $A^{R}_{c}\bar{Q}$ into $\bar{P}_{1}$ and $\bar{P}_{2}$ $O(1)$
+4 Compute $\bar{P}_{2}^{\dagger}$ $O\big{(}24m(n-4p)^{2}+10(n-4p)^{3}\big{)}$
+5 Solve $\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}$ $O(16p^{2}d)$
+6 Compute $\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}$ $O(32mpd-4md)$
+7 Compute $B^{R}_{c}-\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}$ $O(4md)$
+8 Compute $\bar{P}_{2}^{\dagger}(B^{R}_{c}-\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c})$ $O(8m(n-4p)d)$
+9 Compute $X$ $O(2n^{2}d)$
+Total Flop Count $O\big{(}32np^{2}+8mn^{2}-4mn+24m(n-4p)^{2}+10(n-4p)^{3}+16p^{2}d+32mpd-4md+4md+8m(n-4p)d+2n^{2}d\big{)}$
+Table 1: Flop count for the computational steps to find the real solution of the RBLSE problem.
+To analyze the efficiency of the complex solution algorithm, a detailed step-by-step flop count is provided below.
+This analysis emphasizes the additional computations required due to the complex structure.
+Step Description Flop Count
+1 QR decomposition of $\left(C^{C}_{c}\right)^{H}$ $O(32np^{2})$
+2 Compute $A^{C}_{c}\widetilde{Q}$ $O(16mn^{2}-4mn)$
+3 Partition $A^{C}_{c}\widetilde{Q}$ into $\widetilde{P}_{1}$ and $\widetilde{P}_{2}$ $O(1)$
+4 Compute $\widetilde{P}_{2}^{\dagger}$ $O\big{(}48m(n-2p)^{2}+40(n-2p)^{3}\big{)}$
+5 Solve $\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$ $O(16p^{2}d+10pd)$
+6 Compute $\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$ $O(32mpd-4md)$
+7 Compute $B^{C}_{c}-\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$ $O(4md)$
+8 Compute $\widetilde{P}_{2}^{\dagger}(B^{C}_{c}-\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c})$ $O(16m(n-2p)d)$
+9 Compute $X$ $O(8n^{2}d)$
+Total Flop Count $O\big{(}32np^{2}+16mn^{2}-4mn+48m(n-2p)^{2}+40(n-2p)^{3}+16p^{2}d+10pd+32mpd-4md+4md+16m(n-2p)d+8n^{2}d\big{)}$
+Table 2: Flop count for the computational steps to find the complex solution of the RBLSE problem.
 Next, we provide examples to evaluate the effectiveness of the proposed algorithms.
-All computations were performed using a computer equipped with an Intel Core $i7-9700$ processor at $3.00$ GHz and $16$ GB of RAM, running MATLAB $R2021b$ software.
+All computations were performed using a computer equipped with an Intel Core $i9-12900K$ processor at $3200$ MHz and $32$ GB of RAM, running MATLAB $R2024b$ software.
+The following example validates the efficiency of the algorithms by comparing the real and complex solutions in terms of computational accuracy and CPU runtime.
 Example 5.1.
-Let $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}$.
-Let
-$M_{i}$ $=$ $rand(m,n)+rand(m,n)\textit{{i}}\in{\mathbb{C}}^{m\times n},\;\textrm{for}\;i=1,2,$
-$N_{i}$ $=$ $rand(m,d)+rand(m,d)\textit{{i}}\in{\mathbb{C}}^{m\times d},\;\textrm{for}\;i=1,2,$
-$R_{i}$ $=$ $rand(p,n)+rand(p,n)\textit{{i}}\in{\mathbb{C}}^{p\times n},\;\textrm{for}\;i=1,2,$
-$S_{i}$ $=$ $rand(p,d)+rand(p,d)\textit{{i}}\in{\mathbb{C}}^{p\times d},\;\textrm{for}\;i=1,2.$
-Take $m=40t$, $n=6t$, $p=2t$, and $d=3$.
-Here, $t$ is an arbitrary number.
-We apply Algorithm 1 to determine the complex solution of the RBLSE problem.
-Let $X_{CL}$ be the complex solution of the RBLSE problem
-$\min_{X,R_{c}}\left\lVert R_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;AX=B+R_{c},\;CX=D.$
-Let $\epsilon_{1}=\left\lVert AX_{CL}-(B+R_{c})\right\rVert_{F}$ and $\epsilon_{2}=\left\lVert CX_{CL}-D\right\rVert_{F}$.
-In Table 1, we compute $\epsilon_{1}$ and $\epsilon_{2}$ for different values of $t$.
-$t$ $\epsilon_{1}$ $\epsilon_{2}$
-$1$ $8.5131\times 10^{-16}$ $2.0907\times 10^{-15}$
-$3$ $7.0956\times 10^{-16}$ $2.5624\times 10^{-15}$
-$5$ $1.0499\times 10^{-15}$ $3.7683\times 10^{-15}$
-$7$ $1.2804\times 10^{-15}$ $4.5681\times 10^{-15}$
-$9$ $1.4232\times 10^{-15}$ $8.1546\times 10^{-15}$
-Table 1: Computational accuracy of Algorithm 1 for computing the complex solution of the RBLSE problem (11)
-Table 1 shows that the errors $\epsilon_{1}$ and $\epsilon_{2}$ across different values of $t$ are consistently below $10^{-15}$.
-This indicates that Algorithm 1 is highly effective in determining the complex solution for the RBLSE problem (11).
+To evaluate the accuracy and performance of the proposed algorithms for solving the RBLSE problem, we use random data matrices of varying sizes. The problem matrices are defined as follows:
+$A$ $=$ $A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\quad B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$
+$C$ $=$ $C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\quad D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$
+The real matrix components $A_{i},B_{i},C_{i},$ and $D_{i}$ are generated using random values:
+$A_{i}$ $=rand(m,n)\in{\mathbb{R}}^{m\times n},\quad i=0,1,2,3,$
+$B_{i}$ $=rand(m,d)\in{\mathbb{R}}^{m\times d},\quad i=0,1,2,3,$
+$C_{i}$ $=rand(p,n)\in{\mathbb{R}}^{p\times n},\quad i=0,1,2,3,$
+$D_{i}$ $=rand(p,d)\in{\mathbb{R}}^{p\times d},\quad i=0,1,2,3.$
+The matrix dimensions are determined by $m=30t$, $n=10t$, $p=2t$, and $d=2$, where $t$ is a positive integer.
+We apply Algorithms 1 and 2 to compute the real and complex solutions of the RBLSE problem, respectively.
+  1. (a)
+    Let $X_{RL}$ be the real solution of the RBLSE problem:
+    $\min_{X,R_{r}}\|R_{r}\|_{F}\quad\text{subject to}\quad AX=B+R_{r},\quad CX=D.$
+    The accuracy of $X_{RL}$ is measured using the following errors:
+    $\epsilon_{1}=\log_{10}(\|AX_{RL}-(B+R_{r})\|_{F}),\quad\epsilon_{2}=\log_{10}(\|CX_{RL}-D\|_{F}).$
+  2. (b)
+    Let $X_{CL}$ be the complex solution of the RBLSE problem:
+    $\min_{X,R_{c}}\|R_{c}\|_{F}\quad\text{subject to}\quad AX=B+R_{c},\quad CX=D.$
+    The accuracy of $X_{CL}$ is measured using:
+    $\epsilon_{3}=\log_{10}(\|AX_{CL}-(B+R_{c})\|_{F}),\quad\epsilon_{4}=\log_{10}(\|CX_{CL}-D\|_{F}).$
+To thoroughly evaluate the computational performance, the CPU times for each solution are recorded and averaged over $50$ trials to minimize variability.
+The average CPU times (in seconds) for computing $X_{RL}$ and $X_{CL}$ are denoted by $t_{r}$ and $t_{c}$, respectively.
+$t$ $\epsilon_{1}$ $\epsilon_{2}$ $\epsilon_{3}$ $\epsilon_{4}$
+$1$ $-15.1057$ $-15.0513$ $-15.5746$ $-15.2797$
+$3$ $-15.1474$ $-14.5131$ $-15.4137$ $-14.7818$
+$5$ $-14.9570$ $-14.4140$ $-15.3090$ $-14.4512$
+$7$ $-14.8277$ $-14.2473$ $-15.3418$ $-14.5228$
+$9$ $-14.8800$ $-14.0209$ $-15.3164$ $-14.2537$
+Table 3: Accuracy of Algorithms 1 and 2 in computing the real and complex solutions of the RBLSE problem.
+Table 3 demonstrates that the two algorithms achieve high accuracy, with the computational errors consistently below $-14$ for all tested values of $t$.
+This highlights the accuracy of the proposed methods.
+Figure 1: Comparison of CPU time for computing real and complex solution of RBLSE problem
+Figure 1 compares the average CPU times for the two algorithms.
+The results show that the algorithm for computing real solution takes less time compared to the algorithm for computing the complex solution of the RBLSE problem.
+This is consistent with the theoretical complexity of the algorithms.
+The following example compares the theoretical upper bounds with the actual relative forward errors for the real and complex solutions of the RBLSE problem, demonstrating the reliability of the proposed bounds.
 Example 5.2.
-In Example 5.1, we introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively, to examine how these small perturbations affect the complex solution $X_{CL}$ of the RBLSE problem (11).
-These perturbations are measured normwise by the smallest $\epsilon$, as in (21).
-We define the exact relative forward errors for these solutions as $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$.
-Table 2 presents the calculated exact relative forward errors and the corresponding upper bound $U_{CL}$ (calculated using Equation 23) for these solutions across different values of $t$ and $\epsilon$.
-$t$ $\epsilon$ $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$ $U_{CL}$
-$1$ $7.0103\times 10^{-13}$ $1.3055\times 10^{-12}$ $2.7846\times 10^{-11}$
-$7.0580\times 10^{-10}$ $1.3349\times 10^{-9}$ $2.8036\times 10^{-8}$
-$8.9885\times 10^{-7}$ $1.81513\times 10^{-6}$ $3.5704\times 10^{-6}$
-$5$ $7.1245\times 10^{-11}$ $1.1503\times 10^{-10}$ $9.5002\times 10^{-9}$
-$8.9399\times 10^{-9}$ $2.0884\times 10^{-8}$ $1.1921\times 10^{-6}$
-$1.0173\times 10^{-8}$ $1.5357\times 10^{-8}$ $1.3565\times 10^{-6}$
-$9$ $7.1497\times 10^{-12}$ $1.2430\times 10^{-11}$ $1.3370\times 10^{-9}$
-$6.6856\times 10^{-10}$ $1.2308\times 10^{-9}$ $1.2502\times 10^{-7}$
-$1.0212\times 10^{-7}$ $2.5587\times 10^{-7}$ $1.9096\times 10^{-5}$
-Table 2: Comparison of relative forward errors and their upper bounds for the complex solution of a perturbed RBLSE problem (11)
-Table 2 shows that the exact relative forward errors of the complex solution to the RBLSE problem (11) are consistently lower than their respective upper bounds across various values of $t$ and $\epsilon$.
-This verifies the reliability of the derived upper bound $U_{CL}$ for the relative forward error.
+Building on Example 5.1, we now introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively.
+These perturbations are used to examine the sensitivity of the RBLSE problem to small changes in the input data.
+Specifically, we evaluate how these perturbations affect the computed real solution $X_{RL}$ and the complex solution $X_{CL}$.
+The magnitude of these perturbations is measured normwise by the smallest $\epsilon$, which ensures the changes are controlled and small.
+To quantify the impact of the perturbations, we compute the exact relative forward errors for each solution as follows:
+$\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}\quad\text{and}\quad\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}.$
+Here, $\widehat{X}_{RL}$ and $\widehat{X}_{CL}$ denote the computed solutions of the perturbed RBLSE problem, while $X_{RL}$ and $X_{CL}$ are the original solutions.
+In addition to the exact relative forward errors, we compute the corresponding upper bounds $U_{RL}$ and $U_{CL}$ for these errors.
+These bounds are derived using equations (27) and (44), respectively, which provide theoretical estimates of the maximum possible errors.
+Table 4 presents the exact relative forward errors and their corresponding upper bounds for various values of $t$ and $\epsilon$, allowing us to evaluate the relationship between the perturbation magnitude and the stability of the solutions.
+$t$ $\epsilon$ $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$ $U_{RL}$ $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$ $U_{CL}$
+$1$ $1.2957\times 10^{-13}$ $4.5138\times 10^{-13}$ $7.6470\times 10^{-12}$ $3.8724\times 10^{-13}$ $5.6680\times 10^{-12}$
+$1.3279\times 10^{-10}$ $4.3088\times 10^{-10}$ $7.8370\times 10^{-9}$ $1.8961\times 10^{-10}$ $5.8088\times 10^{-9}$
+$1.2770\times 10^{-7}$ $3.0154\times 10^{-7}$ $7.5367\times 10^{-6}$ $2.8896\times 10^{-7}$ $5.5863\times 10^{-6}$
+$5$ $1.2312\times 10^{-11}$ $6.1760\times 10^{-11}$ $4.7674\times 10^{-9}$ $3.1584\times 10^{-11}$ $1.7960\times 10^{-9}$
+$1.5256\times 10^{-9}$ $9.3350\times 10^{-9}$ $5.9071\times 10^{-7}$ $4.8613\times 10^{-9}$ $2.2254\times 10^{-7}$
+$1.4801\times 10^{-8}$ $7.9283\times 10^{-8}$ $5.7313\times 10^{-6}$ $3.3857\times 10^{-8}$ $2.1592\times 10^{-6}$
+$9$ $1.2645\times 10^{-12}$ $6.5592\times 10^{-12}$ $4.5359\times 10^{-10}$ $4.0664\times 10^{-12}$ $2.4556\times 10^{-10}$
+$1.2593\times 10^{-10}$ $6.6261\times 10^{-10}$ $4.5171\times 10^{-8}$ $4.0808\times 10^{-10}$ $2.4454\times 10^{-8}$
+$1.5340\times 10^{-7}$ $7.1997\times 10^{-7}$ $5.5023\times 10^{-5}$ $3.8176\times 10^{-7}$ $2.9788\times 10^{-5}$
+Table 4: Comparison of relative forward errors and their upper bounds for the real and complex solutions of a perturbed RBLSE problem.
+Table 4 illustrates that the exact relative forward errors for the real and complex solutions of the RBLSE problem are consistently smaller than their respective upper bounds $U_{RL}$ and $U_{CL}$, for all considered values of $t$ and $\epsilon$.
+These results validate the accuracy and reliability of the derived upper bounds.
+The following example evaluates the accuracy of the proposed algorithms for computing the real and complex solutions of the RBLSE problem.
 Example 5.3.
-Let $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}$.
-Let
-$A_{i}$ $=$ $randn(m,n)\in{\mathbb{R}}^{m\times n},\;\textrm{for}\;i=0:3,$
-$B_{i}$ $=$ $randn(m,d)\in{\mathbb{R}}^{m\times d},\;\textrm{for}\;i=0:3,$
-$C_{i}$ $=$ $randn(p,n)\in{\mathbb{R}}^{p\times n},\;\textrm{for}\;i=0:3,$
-$D_{i}$ $=$ $randn(p,d)\in{\mathbb{R}}^{p\times d},\;\textrm{for}\;i=0:3.$
-Take $m=30t$, $n=10t$, $p=2t$, and $d=2$.
-Here, $t$ is an arbitrary number.
-We apply Algorithm 2 to determine the real solution of the RBLSE problem.
-Let $X_{RL}$ be the real solution of the RBLSE problem
-$\min_{X,R_{r}}\left\lVert R_{r}\right\rVert_{F}\;\;\textrm{subject to}\;\;AX=B+R_{r},\;CX=D.$
-Let $\epsilon_{1}=\left\lVert AX_{RL}-(B+R_{r})\right\rVert_{F}$ and $\epsilon_{2}=\left\lVert CX_{RL}-D\right\rVert_{F}$.
-In Table 3, we compute $\epsilon_{1}$ and $\epsilon_{2}$ for different values of $t$.
-$t$ $\epsilon_{1}$ $\epsilon_{2}$
-$1$ $2.9063\times 10^{-15}$ $3.0851\times 10^{-15}$
-$3$ $2.5053\times 10^{-15}$ $5.5184\times 10^{-15}$
-$5$ $3.7915\times 10^{-15}$ $1.0949\times 10^{-14}$
-$7$ $3.9204\times 10^{-15}$ $1.3185\times 10^{-14}$
-$9$ $5.2026\times 10^{-15}$ $1.7247\times 10^{-14}$
-Table 3: Computational accuracy of Algorithm 2 for computing the real solution of the RBLSE problem (29)
-Table 3 shows that the errors $\epsilon_{1}$ and $\epsilon_{2}$ across different values of $t$ are consistently below $10^{-14}$.
-This indicates that Algorithm 2 is highly effective in determining the real solution for the RBLSE problem (29).
-Example 5.4.
-In Example 5.3, we introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively, to examine how these small perturbations affect the real solution $X_{RL}$ of the RBLSE problem (29).
-These perturbations are measured normwise by the smallest $\epsilon$, as in (38).
-We define the exact relative forward errors for these solutions as $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$.
-Table 4 presents the computed exact relative forward errors and the corresponding upper bound $U_{RL}$ (calculated using Equation 40) for these solutions across different values of $t$ and $\epsilon$.
-$t$ $\epsilon$ $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$ $U_{RL}$
-$1$ $7.3638\times 10^{-12}$ $2.3814\times 10^{-11}$ $8.0339\times 10^{-10}$
-$7.3677\times 10^{-10}$ $2.1476\times 10^{-9}$ $8.0381\times 10^{-8}$
-$7.9614\times 10^{-8}$ $7.6845\times 10^{-7}$ $8.6858\times 10^{-6}$
-$3$ $7.1742\times 10^{-14}$ $2.6216\times 10^{-13}$ $9.0524\times 10^{-12}$
-$8.1702\times 10^{-11}$ $1.5855\times 10^{-10}$ $1.0309\times 10^{-8}$
-$7.1758\times 10^{-9}$ $2.6438\times 10^{-8}$ $9.0544\times 10^{-7}$
-$5$ $7.2567\times 10^{-11}$ $2.4717\times 10^{-10}$ $1.0430\times 10^{-8}$
-$7.2735\times 10^{-9}$ $2.5130\times 10^{-8}$ $1.0454\times 10^{-6}$
-$8.7538\times 10^{-7}$ $1.8374\times 10^{-6}$ $1.2581\times 10^{-4}$
-Table 4: Comparison of relative forward errors and their upper bounds for the real solution of a perturbed RBLSE problem (29)
-Table 4 shows that the exact relative forward errors of the real solution to the RBLSE problem (29) are consistently lower than their respective upper bounds across various values of $t$ and $\epsilon$.
-This verifies the reliability of the derived upper bound $U_{RL}$ for the relative forward error.
-Example 5.5.
-Let $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times 10}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{3\times 10}$, and $X_{0}=rand(10,3)+rand(10,3)\textit{{i}}\in{\mathbb{C}}^{10\times 3}$.
-Let
-$M_{i}$ $=$ $rand(m,10)+rand(m,10)\textit{{i}}\in{\mathbb{C}}^{m\times 10},\;\textrm{for}\;i=1,2$
-$R_{i}$ $=$ $rand(3,10)+rand(3,10)\textit{{i}}\in{\mathbb{C}}^{3\times 10},\;\textrm{for}\;i=1,2.$
-Take $B=AX_{0}$ and $D=CX_{0}$.
-Thus, $X_{0}$ is the complex solution to the RBLSE problem (11).
-To measure the performance of the proposed technique for the RBLSE problem (11), Algorithm 1 is employed to determine the complex solution $X$.
-Let the error be $\epsilon=\left\lVert X-X_{0}\right\rVert_{F}$.
-Here, $m$ is a variable parameter.
-We evaluate the errors $\epsilon$ for various values of $m$.
-The relationship between the errors $\epsilon$ and $m$ is presented in Table 5.
-$m$ $\epsilon=\|X-X_{0}\|_{F}$
-$100$ $1.3154\times 10^{-14}$
-$200$ $6.2150\times 10^{-15}$
-$300$ $6.5603\times 10^{-15}$
-$400$ $6.1485\times 10^{-15}$
-$500$ $3.2441\times 10^{-15}$
-Table 5: Computational accuracy of Algorithm 1 for computing the complex solution of the RBLSE problem (11)
-Table 5 shows that the error $\epsilon$ between the complex solution derived from Algorithm 1 and the corresponding true solution to the RBLSE problem (11) remains consistently below $10^{-14}$ for various values of $m$.
-This indicates the high accuracy of Algorithm 1 in computing the complex solution to the RBLSE problem (11).
-Example 5.6.
-Let $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times 50}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{10\times 50}$, and $X_{0}=randn(50,30)\in{\mathbb{R}}^{50\times 30}$.
-Let
-$A_{i}$ $=$ $randn(m,50)\in{\mathbb{R}}^{m\times 50},\;\textrm{for}\;i=0:3$
-$C_{i}$ $=$ $randn(10,50)\in{\mathbb{R}}^{10\times 50},\;\textrm{for}\;i=0:3.$
-Take $B=AX_{0}$ and $D=CX_{0}$.
-Clearly, $X_{0}$ is the real solution to the RBLSE problem (29).
-To measure the performance of the proposed technique for the RBLSE problem (29), Algorithm 2 is employed to determine the real solution $X$.
-Define the error as $\epsilon=\left\lVert X-X_{0}\right\rVert_{F}$.
-Here, $m$ is a variable parameter.
-We evaluate the errors $\epsilon$ for various values of $m$.
-The relationship between the errors $\epsilon$ and $m$ is presented in Table 6.
-$m$ $\epsilon=\|X-X_{0}\|_{F}$
-$1000$ $3.8948\times 10^{-14}$
-$2000$ $4.4732\times 10^{-14}$
-$3000$ $4.1257\times 10^{-14}$
-$4000$ $3.7531\times 10^{-14}$
-$5000$ $4.6532\times 10^{-14}$
-Table 6: Computational accuracies of Algorithm 2 for computing the real solution of the RBLSE problem (29)
-Table 6 shows that the error $\epsilon$ between the real solution derived from Algorithm 2 and the corresponding true solution to the RBLSE problem (29) remains consistently below $10^{-14}$ for various values of $m$.
-This indicates the high accuracy of Algorithm 2 in computing the real solution to the RBLSE problem (29).
+We define the problem parameters as follows:
+Let $A\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$ and $C\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$ as defined in Example 5.1.
+For $i=0,1$, let $X_{i}=rand(n,d)\in{\mathbb{R}}^{n\times d}$.
+Using these components, we define:
+$X_{R}=X_{0}\in{\mathbb{R}}^{n\times d},\quad X_{C}=X_{0}+X_{1}\textit{{i}}\in{\mathbb{C}}^{n\times d}$
+Here, the parameters $m$, $p$, $n$, and $d$ remain consistent with Example 5.1 and depend on the variable $t$.
+For the real solution, we set $B=AX_{R}$ and $D=CX_{R}$, which ensures that $X_{R}$ is the exact real solution of the RBLSE problem.
+Algorithm 1 is employed to compute the approximate real solution $\widehat{X}_{R}$.
+The accuracy of the computed solution is evaluated using the Frobenius norm error:
+$\epsilon_{R}=\|X_{R}-\widehat{X}_{R}\|_{F}.$
+For the complex solution, we set $B=AX_{C}$ and $D=CX_{C}$, making $X_{C}$ the exact complex solution of the RBLSE problem.
+Algorithm 2 is used to compute the approximate complex solution $\widehat{X}_{C}$, and the error is defined as:
+$\epsilon_{C}=\|X_{C}-\widehat{X}_{C}\|_{F}.$
+The variable $t$ determines the dimensions of the matrices $A$ and $C$, and the errors $\epsilon_{R}$ and $\epsilon_{C}$ are computed for various values of $t$.
+The results are summarized in Table 5, which shows the relationship between $t$ and the computational accuracy of the algorithms.
+$t$ $\epsilon_{R}=\left\lVert X_{R}-\widehat{X}_{R}\right\rVert_{F}$ $\epsilon_{C}=\left\lVert X_{C}-\widehat{X}_{C}\right\rVert_{F}$
+$1$ $7.0655\times 10^{-15}$ $3.5401\times 10^{-15}$
+$3$ $1.3490\times 10^{-14}$ $1.0711\times 10^{-15}$
+$5$ $2.3275\times 10^{-14}$ $2.2491\times 10^{-14}$
+$7$ $5.4268\times 10^{-14}$ $3.7398\times 10^{-14}$
+$9$ $5.8569\times 10^{-14}$ $5.2257\times 10^{-14}$
+Table 5: Computational accuracy of Algorithms 1 and 2 for computing the real and complex solutions of the RBLSE problem.
+Table 5 demonstrates that the errors $\epsilon_{R}$ and $\epsilon_{C}$ are consistently below $10^{-14}$ for all tested values of $t$.
+These results confirm the high accuracy of the proposed algorithms for computing the real and complex solutions of the RBLSE problem, showcasing their effectiveness and precision across varying problem sizes.
 6 Conclusions
-We have presented an algebraic method for solving the RBLSE problem by transforming it into complex and real LSE problems.
-This transformation leverages the complex and real representations of reduced biquaternion matrices, enabling efficient computation of both complex and real solutions to the RBLSE problem.
-Additionally, we have calculated the upper bound for the relative forward error associated with these solutions, demonstrating the robustness and accuracy of our method in addressing the RBLSE problem.
+We have presented an algebraic method for solving the RBLSE problem by transforming it into real and complex LSE problems.
+This transformation leverages the real and complex representations of reduced biquaternion matrices, enabling efficient computation of real and complex solutions to the RBLSE problem.
+Additionally, we have calculated the upper bound for the relative forward error associated with these solutions, demonstrating the accuracy and reliability of our method in addressing the RBLSE problem.
+Future research will focus on exploring their applications in image and signal processing.
 7 Declarations
@@
 The first author played a crucial role in overseeing the project and providing valuable guidance.
-Neha Bhadala made significant contributions at various stages of the research, including conceptualization, methodology design, software implementation, validation, and formal analysis.
-Additionally, Neha Bhadala took the lead in drafting the original manuscript and actively participated in the subsequent writing and editing phases.
+Neha Bhadala made significant contributions at various stages of the research, including conceptu-alization, methodology design, software implementation, validation, and formal analysis.
+Additio-nally, Neha Bhadala took the lead in drafting the original manuscript and actively participated in the subsequent writing and editing phases.
+Conflict of interest
+This work does not have any conflicts of interest.
 Acknowledgement